### v6

# DualDynamics: Synergizing Implicit and Explicit Methods for Robust Irregular Time Series Analysis

Abstract
Real-world time series analysis faces significant challenges when dealing with irregular and incomplete data. While Neural Differential Equation (NDE) based methods have shown promise, they struggle with limited expressiveness, scalability issues, and stability concerns. Conversely, Neural Flows offer stability but falter with irregular data. We introduce DualDynamics, a novel framework that synergistically combines NDE-based method and Neural Flow-based method. This approach enhances expressive power while balancing computational demands, addressing critical limitations of existing techniques. We demonstrate DualDynamics’ effectiveness across diverse tasks: classification of robustness to dataset shift, irregularly-sampled series analysis, interpolation of missing data, and forecasting with partial observations. Our results show consistent outperformance over state-of-the-art methods, indicating DualDynamics’ potential to advance irregular time series analysis significantly.
Codehttps://github.com/yongkyung-oh/DualDynamics
Extended versionhttps://arxiv.org/abs/2401.04979
Introduction
Effectively modeling time series data is a cornerstone in machine learning, supporting numerous applications across diverse sectors. This significance has driven substantial research endeavors aimed at developing novel methodologies capable of accurately capturing the complexities of continuous latent processes (Chen et al. 2018; Kidger et al. 2020; Kidger 2022; Oh, Lim, and Kim 2024).
Recent advancements have led to two distinct approaches: implicit methods, such as Neural Differential Equation (NDE)-based techniques, and explicit methods, such as Neural Flows. Neural Ordinary Differential Equations (Neural ODEs) (Chen et al. 2018), Neural Controlled Differential Equations (Neural CDEs) (Kidger et al. 2020) and Neural Stochastic Differential Equations (Neural SDEs) (Oh, Lim, and Kim 2024) have emerged as prominent implicit methods, recognized for their ability to learn continuous-time dynamics and underlying temporal structures. These approaches directly learn continuous latent representations based on vector fields parameterized by neural networks. However, NDE-based methods face challenges including limited expressive power (Dupont, Doucet, and Teh 2019; Massaroli et al. 2020b; Chalvidal et al. 2021; Kidger 2022) and scalability issues when analyzing irregular or complex time series data (Norcliffe et al. 2020; Morrill et al. 2021, 2022; Irie, Faccio, and Schmidhuber 2022; Pal et al. 2023). These limitations are attributed to data complexity, sequence length variations, and the stability constraints of numerical solvers (He and Semnani 2023; Westny et al. 2023).
In parallel, explicit methods have been developed, focusing on directly mapping solution curves using neural networks (Lu et al. 2018; Sonoda and Murata 2019; Massaroli et al. 2020a; Biloš et al. 2021). These approaches offer invertible solutions and enhanced stability through the change of variables formula (Kobyzev, Prince, and Brubaker 2020; Papamakarios et al. 2021). However, they struggle with irregularly-sampled time series and are sensitive to initial state and corresponding initial value problem.
To address the limitations of both implicit and explicit methods, we introduce DualDynamics, a novel framework that synergistically combines NDE-based model and Neural Flow-based model. This approach aims to leverage the strengths of both paradigms: the flexibility of implicit methods in handling irregular data and the stability and computational efficiency of explicit methods. By integrating these complementary approaches, DualDynamics seeks to enhance expressive power, improve scalability, and maintain stability in modeling complex, irregular time series data.
Related Works
The landscape of neural differential equations for time series modeling has been dominated by two distinct approaches: implicit methods, exemplified by Neural ODEs, Neural CDEs and Neural SDEs, and explicit methods, such as Neural Flows. These approaches fundamentally differ in how they represent and solve the underlying temporal dynamics.
Implicit Methods for Continuous Dynamics
Implicit methods in deep learning for time series modeling solve differential equations numerically, allowing for adaptive computation and flexible handling of irregular data. These methods have gained significant attention due to their ability to model complex dynamics.
Let ${\bm{x}}=(x_{0},x_{1},\ldots,x_{n})$ be a vector of original irregularly-sampled observations and ${\bm{z}}(t)$ denote a latent state at time $t$. Continuous latent dynamics ${\bm{z}}(t)$ can be achieved by integration of ${\textnormal{d}}{\bm{z}}(t)$ over time in $[0,t]$. Neural Differential Equations can be expressed as:
[…]
$+g(s,{\bm{z}}(s);\theta_{g})\,{\textnormal{d}}W(s)$ (3)
with initial condition ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$, where $h:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{z}}$ is an affine function. $f$ and $g$ are neural networks parameterized by $\theta_{f}$ and $\theta_{g}$, respectively. $X(t)$ is a control path for Neural CDEs, and $W(t)$ is a Wiener process (or Brownian motion) for Neural SDEs.
Neural ODEs (Chen et al. 2018) model deterministic dynamics but may struggle with expressiveness and stability for complex trajectories (Dupont, Doucet, and Teh 2019; Massaroli et al. 2020b). Furthermore, Neural CDEs (Kidger et al. 2020) handle irregular time series effectively but can be computationally intensive for long sequences (Morrill et al. 2021). Therefore, recent works introduced extended architecture using Neural CDE. On the other hand, Neural SDEs (Li et al. 2020) incorporate uncertainty modeling but face challenges in training and inference due to their stochastic nature (Jia and Benson 2019; Oh, Lim, and Kim 2024). These methods offer varying approaches to modeling continuous-time dynamics in time series data, each with its own strengths and limitations in handling complexity, irregularity, and uncertainty.
Explicit Methods for Continuous Dynamics
Explicit methods directly model the solution of differential equations, often providing faster computation and improved stability, albeit potentially at the cost of some flexibility. Neural Flows, introduced by Biloš et al. (2021), present a novel approach to directly model the trajectory of the latent state ${\bm{z}}(t)$, in contrast to Neural ODEs which represent the rate of change of ${\bm{z}}(t)$ over time in an integration form. Specifically, ${\bm{z}}(t)$ is expressed as:
${\bm{z}}(t)$ $={\mathcal{F}}(t,{\bm{z}}(0);\theta_{\mathcal{F}}),$ (4)
with the initial value ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$, where ${\mathcal{F}}$ is a neural network parameterized by $\theta_{\mathcal{F}}$. Note that while $\mathcal{F}$ represents the solution to the initial value problem of Neural ODEs, as described in Eq. (1), the formulation of Neural Flows in Eq. (4) does not require the use of an ODE solver. However, the requirement for the Neural Flow ${\mathcal{F}}(\cdot,{\bm{z}}(0))$ to be invertible restricts the type of neural network architectures, which may not be suitable if the initial value is unknown or unstable to estimate. In practice, determining the exact flow is challenging since many real-world problems do not have analytical solutions, which can lead to training instability.
Methodology
Proposed Framework
To illustrate our DualDynamics framework, we utilize Neural CDEs as the primary example in this section. Neural CDEs serve as an effective representation of the implicit component in our dual approach, offering a robust foundation for modeling irregular time series data.
Figure 1 shows the conceptual overview of the proposed architecture using Neural CDE (implicit) and Flow-based model (explicit) for classification task.
Figure 1: Overview of the proposed DualDynamics
The proposed method combines the stability of flow models with NDE-based model to achieve a better representation of the irregular time series. The flow model, known for its capability to smoothly transform complex distributions, offers inherent stability when dealing with time-varying dynamics. To achieve stability in the latent representation ${\bm{z}}(t)$, we introduce a secondary latent space, denoted by $\hat{{\bm{z}}}(t)$. This secondary latent space is inspired by the principles of Neural Flow and is designed to produce a regularized representation of the original latent variable. From Eq. (2), the latent representation ${\bm{z}}(t)$ can be derived from an initial vector ${\bm{z}}(0)$, which is transformed from the raw input $x_{0}$. Thus, the proposed dual latent variable evolves in time as:
${\bm{z}}(t)$ $={\bm{z}}(0)+\int_{0}^{t}f(s,{\bm{z}}(s);\theta_{f})\,{\textnormal{d}}X(s),$ (5)
where the initial value of the first latent representation ${\bm{z}}(0)=h(x_{0};\theta_{h})$, and $h:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{z}}$. As stated in Chen et al. (2018) and Kidger et al. (2020), we focus on the time-evolving state ${\bm{z}}$ from the linear mapping from ${\bm{x}}$ using $h$. Eq. (5) can be rewritten as follows:
${\bm{z}}(t)$ $={\bm{z}}(0)+\int_{0}^{t}f^{*}(s;\theta_{f},X)\,{\textnormal{d}}s,$ (6)
where $f^{*}(s;\theta_{f},X):=f(s,{\bm{z}}(s);\theta_{f})\frac{{\textnormal{d}}X(s)}{{\textnormal{d}}s}$. Recall that $X$ is generated from ${\bm{x}}$. Here, we implement $X$ for ‘Hermite Cubic splines with a backward difference’, which is smooth and unique, as discussed by Morrill et al. (2022). Through this part, we find ${\bm{z}}\in\mathbb{R}^{d_{z}}$, and then plug it into the following flow model ${\mathcal{G}}$ as:
$\hat{{\bm{z}}}(t)$ $={\mathcal{G}}(t,\hat{{\bm{z}}}(0);\theta_{\mathcal{G}}),$ (7)
with $\hat{{\bm{z}}}(0)=k({\bm{z}};\theta_{k})$, and $k:\mathbb{R}^{d_{z}\times T}\rightarrow\mathbb{R}^{d_{z}}$. In other words, $\hat{{\bm{z}}}(0)$ is determined by ${\bm{z}}(t)$ for $t\in[0,T]$.
Here, ${\mathcal{G}}$ satisfies the ODE defined in Eq. (6), i.e., $\frac{\mathrm{d}{\mathcal{G}}(t,\hat{{\bm{z}}}(0))}{\mathrm{d}t}=f^{*}(t;\theta_{f},X)$ and ${\mathcal{G}}(\cdot,\hat{{\bm{z}}}(0))$ is invertible. In summary, after obtaining the primary latent space ${\bm{z}}(t)$, we compute ${\mathcal{G}}$ with the initial value $\hat{{\bm{z}}}(0)$. This model generates the secondary latent state $\hat{{\bm{z}}}(t)$ without the need to directly determine its derivative $\frac{{\textnormal{d}}\hat{{\bm{z}}}(t)}{{\textnormal{d}}t}$.
Expressive Power of Flow Model ${\mathcal{G}}$
Flow-based model, which is invertible and differentiable (known as ‘diffeomorphisms’), has the capability of representing any distribution. Given the Neural Flow ${\mathcal{G}}$, there exists a unique input $\hat{{\bm{z}}}(0)$ such that, $\hat{{\bm{z}}}(0)={\mathcal{G}}^{-1}(t,\hat{{\bm{z}}}(t);\theta_{\mathcal{G}})$. This invertibility ensures a one-to-one mapping between the initial and transformed states, preserving information content. The change of variables formula for probability densities leads to:
$p_{\hat{{\bm{z}}}(t)}(\hat{{\bm{z}}}(t))=p_{\hat{{\bm{z}}}(0)}(\hat{{\bm{z}}}(0))\left|\det\bm{J}_{{\mathcal{G}}}(\hat{{\bm{z}}}(0))\right|^{-1},$ (8)
where $\bm{J}_{{\mathcal{G}}}(\hat{{\bm{z}}}(0))$ is the Jacobian matrix of ${\mathcal{G}}$ at $\hat{{\bm{z}}}(0)$, and $\left|\det\bm{J}_{{\mathcal{G}}}(\hat{{\bm{z}}}(0))\right|$ is its determinant.
Suppose $\hat{{\bm{z}}}(0)\sim\mathcal{U}([0,1]^{d_{z}})$, meaning each component of $\hat{{\bm{z}}}(0)$ is uniformly distributed between 0 and 1. The probability density function of $\hat{{\bm{z}}}(0)$ in this multi-dimensional space is constant, $p_{\hat{{\bm{z}}}(0)}(y)=1$ if $y\in[0,1]^{d_{z}}$. Then, Eq. (8) can be simplified to:
$p_{\hat{{\bm{z}}}(t)}(\hat{{\bm{z}}}(t))=\left|\det\bm{J}_{{\mathcal{G}}}(\hat{{\bm{z}}}(0))\right|^{-1}.$ (9)
The determinant of the Jacobian $\left|\det\bm{J}_{{\mathcal{G}}}(\hat{{\bm{z}}}(0))\right|$ accounts for the ‘stretching’ or ‘squeezing’ of the volume elements during the transformation. By adjusting $\theta_{\mathcal{G}}$, the model can control how much each part of the space $[0,1]^{d_{z}}$ expands, contracts or projects, allowing ${\mathcal{G}}$ to mold the uniform distribution into any desired complex distribution $p_{\hat{{\bm{z}}}(t)}$.
Properties of the Proposed Method
Preservation of Probability Density
It is crucial for ensuring that the total probability across transformations remains constant, enabling accurate modeling of dynamic systems’ evolution without loss or gain of information (Li and Chen 2008; Dinh, Sohl-Dickstein, and Bengio 2016). For each fixed $t$, $p_{{\bm{z}}(t)}$ and $p_{\hat{{\bm{z}}}(t)}$ be the probability density functions of ${\bm{z}}(t)$ and $\hat{{\bm{z}}}(t)$, respectively. From the instantaneous change of variables formula (Chen et al. 2018), we observe that:
$\frac{\partial\log p_{{\bm{z}}(t)}({\bm{z}}(t))}{\partial t}$ $=-\text{Tr}\left(\bm{J}_{f^{*}}({\bm{z}}(t))\right),$ (10)
$\frac{\partial\log p_{\hat{{\bm{z}}}(t)}(\hat{{\bm{z}}}(t))}{\partial t}$ $=-\text{Tr}\left(\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}t}}(\hat{{\bm{z}}}(t))\right),$ (11)
where $Tr(A)$ is the trace of matrix $A$ and $\bm{J}_{f^{*}}$ is the Jacobian matrix of $f^{*}$. $\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}t}}$ is the Jacobian matrix of $\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}t}$ and we have used $\frac{\mathrm{d}{\mathcal{G}}(t,\hat{{\bm{z}}}(0))}{\mathrm{d}t}=f^{*}(t;\theta_{f},X)$ for all $t$. Therefore, the total changes in log-density for ${\bm{z}}$ and $\hat{{\bm{z}}}$ from time 00 to $t$ are given by:
$\log p_{{\bm{z}}(t)}({\bm{z}}(t))-\log p_{{\bm{z}}(0)}({\bm{z}}(0))$
$\qquad=-\int_{0}^{t}\text{Tr}\left(\bm{J}_{f^{*}}({\bm{z}}(s))\right)\,{\textnormal{d}}s,$ (12)
$\log p_{\hat{{\bm{z}}}(t)}(\hat{{\bm{z}}}(t))-\log p_{\hat{{\bm{z}}}(0)}(\hat{{\bm{z}}}(0))$
$\qquad=-\int_{0}^{t}\text{Tr}\left(\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}s}}(\hat{{\bm{z}}}(s))\right)\,{\textnormal{d}}s,$ (13)
$\qquad=-\int_{0}^{t}\text{Tr}\left(\bm{J}_{f^{*}}(\hat{{\bm{z}}}(s))\right)\,{\textnormal{d}}s,$ (14)
Eq. (12) and Eq. (14) imply that the probability content of corresponding regions in the primary latent space ${\bm{z}}$ and the second latent space $\hat{\bm{z}}$ must be the same when the Neural Flow ${\mathcal{G}}$ is perfectly estimated (Kingma and Dhariwal 2018; Onken et al. 2021).
Computational Advantage of Flow Model ${\mathcal{G}}$
We are interested in computing the trace of $\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}s}}$ in Eq. (13), which is a key quantity to understand the change in log-density. We use Hutchinson’s trace estimator111 Hutchinson’s trace estimator states that for a square matrix $A$, the trace of $A$ can be estimated as: $\text{Tr}(A)\approx\mathbb{E}_{\bm{v}}[\bm{v}^{\top}A\bm{v}],$ where $\bm{v}$ is a random vector with each element drawn independently from a distribution with zero mean and unit variance. (Hutchinson 1989) used to estimate the trace of a matrix efficiently, especially in situations where the matrix is large and computing the trace directly is computationally expensive [^2] Hutchinson’s trace estimator significantly reduces computational complexity from $\mathcal{O}({d_{z}}^{2})$ for direct trace computation to $\mathcal{O}(d_{z})$ for matrix-vector multiplications. (Adams et al. 2018; Han, Malioutov, and Shin 2015; Grathwohl et al. 2018).
Applying Hutchinson’s trace estimator to $\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}s}}$, we have:
$\text{Tr}\left(\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}t}}(\hat{{\bm{z}}}(t))\right)\approx\mathbb{E}_{\bm{v}}\left[\bm{v}^{\top}\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}t}}(\hat{{\bm{z}}}(t))\bm{v}\right].$ (15)
Then, Eq. (13) can be rewritten as follows:
$\log p_{\hat{{\bm{z}}}(t)}(\hat{{\bm{z}}}(t))-\log p_{\hat{{\bm{z}}}(0)}(\hat{{\bm{z}}}(0))$
$\qquad\approx-\mathbb{E}_{\bm{v}}\left[\int_{0}^{t}\bm{v}^{\top}\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}t}}(\hat{{\bm{z}}}(t))\bm{v}\,{\textnormal{d}}s\right]$
$\qquad=-\mathbb{E}_{\bm{v}}\left[\int_{0}^{t}\bm{v}^{\top}\frac{\partial\bm{J}_{{\mathcal{G}}}(\hat{{\bm{z}}}(t))}{\partial t}\bm{v}\,{\textnormal{d}}s\right].$ (16)
where the last inequality is obtained by changing the order of partial derivatives. Incorporating the flow model ${\mathcal{G}}$ into the original NDE-based model enhances its expressive power without significantly increasing the computational burden, as opposed to the addition of an extra NDE formulation.
Parameter Optimization
To optimize parameters involved in the proposed method, we employ adjoint-based backpropagation (Chen et al. 2018; Kidger et al. 2020; Kidger 2022; Oh, Lim, and Kim 2024). This method allows for efficient gradient computation, ensuring that the information from one step influences and refines the subsequent step, leading to a more unified optimization process. The adjoint method is a powerful technique often used to calculate the gradients of systems governed by differential equations. The basic idea is to recast the derivative computation problem as the adjoint equation (Pontryagin 2018; Pollini, Lavan, and Amir 2018).
Consider an objective function $\mathcal{L}({\bm{z}},\hat{{\bm{z}}}(T);\theta_{f},\theta_{{\mathcal{G}}})$ that depends on the final value of the second latent state, $\hat{\bm{z}}(T)$. Then, one can define the adjoint state $\bm{\lambda}_{{\bm{z}}}(t)$ and $\bm{\lambda}_{\hat{{\bm{z}}}}(t)$ as the gradient of the loss function with respect to the state ${\bm{z}}(t)$ and $\hat{{\bm{z}}}(t)$:
$\bm{\lambda}_{{\bm{z}}}(t)=\frac{\partial\mathcal{L}}{\partial{\bm{z}}(t)},\qquad\bm{\lambda}_{\hat{{\bm{z}}}}(t)=\frac{\partial\mathcal{L}}{\partial\hat{{\bm{z}}}(t)}.$ (17)
To estimate the model parameters $\theta_{f}$ and $\theta_{\mathcal{G}}$, we compute the gradients of the loss function with respect to these parameters simultaneously:
$\frac{\partial\mathcal{L}}{\partial\theta_{f}}$ $=\int_{0}^{T}\bm{\lambda}_{{\bm{z}}}(t)^{\top}\frac{\partial f(t,{\bm{z}}(t);\theta_{f})}{\partial\theta_{f}}\,\frac{{\textnormal{d}}X(t)}{{\textnormal{d}}t}{\textnormal{d}}t,$ (18)
$\frac{\partial\mathcal{L}}{\partial\theta_{\mathcal{G}}}$ $=\bm{\lambda}_{\hat{{\bm{z}}}}(T)^{\top}\frac{\partial{\mathcal{G}}(T,\hat{{\bm{z}}}(0);\theta_{{\mathcal{G}}})}{\partial\theta_{\mathcal{G}}}.$ (19)
Note that $\bm{\lambda}_{{\bm{z}}}(t)$ can be derived by integration of adjoint dynamics backward in time from $T$ to 0, which is the continuous-time analog to the backpropagation algorithm (Rumelhart, Hinton, and Williams 1986; Kidger, Chen, and Lyons 2021). On the other hand, $\bm{\lambda}_{\hat{{\bm{z}}}}(t)$ is derived by backpropagation algorithm directly.
The DualDynamics framework integrates implicit and explicit components into a unified architecture, enabling concurrent optimization of both elements. This joint learning process ensures that the implicit Neural Differential Equation and the explicit Neural Flow components evolve synergistically, rather than sequentially, thereby capitalizing on their complementary strengths in a single, cohesive optimization procedure.
Experiments
All experiments were performed using a server on Ubuntu 22.04 LTS, equipped with an Intel(R) Xeon(R) Gold 6242 CPU and a cluster of NVIDIA A100 40GB GPUs. The source code for our experiments and datasets can be accessed at https://github.com/yongkyung-oh/DualDynamics. Please refer the supplementary material for the experiment settings and implementation details.
Proposed Method
The implicit component in our framework primarily utilizes Neural CDE, with Neural ODE and Neural SDE explored as alternatives in our ablation study. For the explicit component, we evaluated three distinct flow models: ResNet Flow, GRU Flow, and Coupling Flow. These flow architectures, as outlined by Biloš et al. (2021), were incorporated into our proposed framework. To assess the impact of using flow models, we also implemented a conventional multilayer perceptron (MLP) as a baseline for comparison. In practice, optimal flow model is selected by hyperparameter tuning stage.
Benchmark Methods
In experiments, we employed a variety of benchmark models for classification and forecasting, incorporating conventional RNN (Rumelhart, Hinton, and Williams 1986; Medsker and Jain 1999), LSTM (Hochreiter and Schmidhuber 1997), and GRU (Chung et al. 2014), alongside their notable variants, including GRU-$\Delta t$ (Choi et al. 2016), and GRU-D (Che et al. 2018). Also, we used attention-based method including transformer (Vaswani et al. 2017), Multi-Time Attention Networks (MTAN) (Shukla and Marlin 2021), and Multi-Integration Attention Module (MIAM) (Lee et al. 2022). ODE-based models like, GRU-ODE (De Brouwer et al. 2019), ODE-RNN (Rubanova, Chen, and Duvenaud 2019) ODE-LSTM (Lechner and Hasani 2020), Latent-ODE (Rubanova, Chen, and Duvenaud 2019), Augmented-ODE (Dupont, Doucet, and Teh 2019), and Attentive co-evolving neural ordinary differential equations (ACE-NODE) (Jhin et al. 2021) are considerd. Furthermore, Neural CDE (Kidger et al. 2020), and Neural Rough Differential Equation (Neural RDE) (Morrill et al. 2021), were included in our comparative study. Additionally, recent advances in Neural CDE are considered, including Attentive Neural Controlled Differential Equation (ANCDE) (Jhin et al. 2024), EXtrapolation and InTerpolation-based model (EXIT) (Jhin et al. 2022), and LEArnable Path-based model (LEAP) (Jhin et al. 2023). Lastly, variations of Neural Flow (Biloš et al. 2021) are also included.
In the interpolation task, encoder-decoder architecture is implemented by Variational Auto-Encoder (VAE) scheme with the evidence lower bound (ELBO). RNN-VAE (Chen et al. 2018), L-ODE-RNN (Chen et al. 2018), L-ODE-ODE (Rubanova, Chen, and Duvenaud 2019), and mTAND-Full (MTAN encoder–MTAN decoder model) (Shukla and Marlin 2021), were used as benchmark methods, based on the suggestion of Shukla and Marlin (2021). In our model, the proposed method is utilized for the encoder component, while the conventional RNN is employed as the decoder.
Methods Regular 30% Missing 50% Missing 70% Missing Average
Accuracy Rank Accuracy Rank Accuracy Rank Accuracy Rank Accuracy Rank
RNN 0.560 (0.072) 11.9 0.484 (0.075) 14.7 0.471 (0.082) 14.0 0.453 (0.068) 14.4 0.492 (0.074) 13.8
LSTM 0.588 (0.067) 11.0 0.552 (0.075) 10.2 0.516 (0.073) 11.4 0.505 (0.067) 11.3 0.540 (0.071) 11.0
GRU 0.674 (0.080) 7.3 0.639 (0.065) 8.4 0.611 (0.076) 9.0 0.606 (0.088) 8.6 0.633 (0.077) 8.3
GRU-$\Delta t$ 0.629 (0.065) 9.9 0.636 (0.069) 8.2 0.651 (0.068) 7.2 0.649 (0.074) 7.9 0.641 (0.069) 8.3
GRU-D 0.593 (0.088) 10.9 0.579 (0.087) 10.6 0.580 (0.075) 10.6 0.599 (0.062) 10.1 0.588 (0.078) 10.6
Transformer 0.720 (0.063) 6.8 0.663 (0.066) 8.0 0.669 (0.075) 7.4 0.643 (0.079) 8.2 0.674 (0.071) 7.6
MTAN 0.654 (0.088) 9.4 0.634 (0.098) 7.4 0.631 (0.089) 6.8 0.642 (0.073) 6.8 0.640 (0.087) 7.6
MIAM 0.573 (0.079) 11.9 0.585 (0.086) 10.0 0.572 (0.071) 9.9 0.544 (0.063) 11.1 0.568 (0.075) 10.7
GRU-ODE 0.663 (0.072) 7.6 0.661 (0.069) 7.4 0.664 (0.069) 6.8 0.659 (0.081) 6.9 0.662 (0.073) 7.2
ODE-RNN 0.652 (0.085) 7.3 0.632 (0.076) 7.8 0.626 (0.086) 7.8 0.653 (0.059) 6.5 0.641 (0.076) 7.4
ODE-LSTM 0.566 (0.074) 11.4 0.518 (0.069) 12.7 0.501 (0.068) 13.3 0.474 (0.068) 13.4 0.515 (0.070) 12.7
Neural CDE 0.681 (0.073) 7.6 0.672 (0.068) 7.4 0.661 (0.070) 7.3 0.652 (0.091) 7.3 0.667 (0.075) 7.4
Neural RDE 0.649 (0.082) 8.6 0.648 (0.071) 7.4 0.633 (0.078) 8.2 0.607 (0.079) 8.5 0.634 (0.078) 8.2
ANCDE 0.662 (0.083) 7.9 0.661 (0.083) 7.1 0.639 (0.080) 8.2 0.631 (0.073) 7.3 0.649 (0.080) 7.6
EXIT 0.595 (0.087) 10.1 0.580 (0.088) 10.5 0.578 (0.086) 10.1 0.564 (0.072) 10.4 0.579 (0.083) 10.3
LEAP 0.490 (0.062) 14.4 0.459 (0.070) 15.1 0.466 (0.074) 13.6 0.451 (0.074) 13.7 0.466 (0.070) 14.2
Neural Flow 0.545 (0.059) 11.9 0.485 (0.067) 13.4 0.455 (0.058) 14.3 0.438 (0.054) 14.1 0.481 (0.059) 13.4
Proposed method 0.724 (0.090) 5.1 0.720 (0.088) 4.9 0.691 (0.091) 5.1 0.697 (0.098) 4.6 0.708 (0.092) 4.9
Table 1: Average classification performance on 18 datasets under regular and three missing rates (Values in parentheses show the average of 18 standard deviations. The top-ranking and runner-up methods are emphasized.)
Robustness to Dataset Shift
Descriptive Example.
To provide comprehensive insights, we conducted a descriptive analysis of model performances using the ‘BasicMotions’ dataset as a case study. We focused on five key methods: a standard RNN, GRU-ODE, ODE-LSTM, Neural CDE, Neural Flow, and our proposed method. Our investigation covered four different settings, including regular (0% missingness), and three missing rates including 30%, 50%, and 70%, respectively.
Figure 2: Classification performance comparison on the ‘BasicMotions’ dataset with different scenarios
Figure 2 illustrates the classification performance trend of each model concerning varying missing data rates. While the conventional RNN’s performance decreases as the missing rate increases, GRU-ODE and ODE-LSTM outperform conventional RNNs but become less stability with increasing missing data. Neural CDE demonstrates superior performance, whereas Neural Flow shows performance worse than that of the conventional RNN.
(a) Test loss without missingness (regular)
(b) Test loss with 50% missing (irregular)
Figure 3: Stability comparison on the ‘BasicMotions‘ dataset, monitored over 100 epochs without early stopping
Our proposed method consistently demonstrates remarkable performance regardless of missing data rates. Further exploration is presented in Figure 3 (a) and (b), illustrating the test loss dynamics of the ‘BasicMotions’ dataset, both with and without missingness. Despite its simplicity, Neural Flow shows limited convergence in both cases, whereas our proposed method achieves stability, outperforming both conventional approaches and Neural CDE. Furthermore, we included detailed results and computational time analysis in the supplementary material.
Configuration Accuracy Loss
Proposed method (Neural CDE) 0.708 (0.092) 0.686 (0.168)
[…]
– Conventional MLP 0.582 (0.087) 0.862 (0.148)
– ResNet Flow 0.651 (0.091) 0.777 (0.155)
– GRU Flow 0.660 (0.079) 0.768 (0.172)
– Coupling Flow 0.669 (0.097) 0.747 (0.186)
Proposed method (Neural ODE) 0.535 (0.066) 0.968 (0.072)
[…]
– Conventional MLP 0.504 (0.066) 0.980 (0.076)
– ResNet Flow 0.510 (0.060) 0.978 (0.077)
– GRU Flow 0.523 (0.071) 0.981 (0.080)
– Coupling Flow 0.523 (0.064) 0.970 (0.070)
Proposed method (Neural SDE) 0.538 (0.066) 0.968 (0.071)
[…]
– Conventional MLP 0.506 (0.069) 0.981 (0.077)
– ResNet Flow 0.520 (0.064) 0.969 (0.073)
– GRU Flow 0.521 (0.068) 0.981 (0.073)
– Coupling Flow 0.519 (0.068) 0.975 (0.079)
Table 2: Ablation study of the proposed method
Methods Test MSE ×$\times 10^{-3}$
50% 60% 70% 80% 90%
RNN-VAE 13.418 ± 0.008 12.594 ± 0.004 11.887 ± 0.005 11.133 ± 0.007 11.470 ± 0.006
L-ODE-RNN 8.132 ± 0.020 8.140 ± 0.018 8.171 ± 0.030 8.143 ± 0.025 8.402 ± 0.022
L-ODE-ODE 6.721 ± 0.109 6.816 ± 0.045 6.798 ± 0.143 6.850 ± 0.066 7.142 ± 0.066
mTAND-Full 4.139 ± 0.029 4.018 ± 0.048 4.157 ± 0.053 4.410 ± 0.149 4.798 ± 0.036
Proposed method 3.631 ± 0.049 3.659 ± 0.028 3.463 ± 0.032 3.224 ± 0.044 3.114 ± 0.050
Table 3: Performance of interpolation relative to observed percentage on the PhysioNet Mortality dataset
Comprehensive Analysis.
Based on our initial findings, we extended our experiments to 18 public benchmark datasets from the the University of East Anglia (UEA) and the University of California Riverside (UCR) Time Series Classification Repository (Bagnall et al. 2018; Dau et al. 2019), encompassing Motion & Human Activity Recognition (HAR), Electrocardiogram (ECG) & Electroencephalogram (EEG), and Sensor domains. Following Kidger et al. (2020) and Oh, Lim, and Kim (2024), we generated subsets with 30%, 50%, and 70% missingness rates, creating four distinct scenarios. Data was split 70:15:15 for training, validation, and testing. We employed five repetition, evaluating mean performance and ranking across iterations.
In the results presented in Table 1, our method distinctly stands out, demonstrating superiority when compared with other contemporary techniques. Considering the volatile nature of accuracy scores across diverse datasets, we prioritize rank statistics as a more consistent metric for comparative analysis. Across the four missing rate scenarios considered, our proposed method consistently achieves top-tier accuracy, confirming its robustness and effectiveness.
We conducted an in-depth comparative analysis of the classification efficacy and cross-entropy loss across various flow model designs, as detailed in Table 2. For implicit component, Neural CDE outperformed compared to Neural ODE and Neural SDE. The proposed method selected the flow configuration from ResNet flow, GRU flow, and Coupling flow during hyperparameter tuning. For the comparison, we employed conventional MLP, which notably failed to meet specific flow criteria. Empirical evaluations indicate that the performance of conventional MLP is worse compared to the flow-based approach. This observation confirms the superiority of our proposed method.
Interpolation of Missing Data
The interpolation experiment utilized the 2012 PhysioNet Mortality dataset (Silva et al. 2012), comprising 37 variables from ICU records over 48 hours post-admission. Following Rubanova, Chen, and Duvenaud (2019), timestamps were adjusted to the nearest minute, yielding up to 2880 intervals per series. We adhered to Shukla and Marlin (2021)’s methodology, varying observation frequency from 50% to 90% to predict remaining samples using 8000 instances.
In our interpolation experiments, we followed the experimental protocol advocated by Shukla and Marlin (2021). This strategy involved generating interpolated values based on a selected subset of data points to predict values for the unobserved time intervals. The interpolation spanned an observational range from 50% to 90% of the total data points. During the test phase, models were designed to interpolate values for the remaining intervals in the test set. The efficacy of the models was quantitatively assessed using Mean Squared Error (MSE), supported by five cross-validations.
Table 3 demonstrates that the interpolation performance of our proposed methods is superior compared to the benchmark models across all three flow configurations. The flow configuration achieved lower MSE errors compared to the non-flow neural network, a consistent and robust observation across various levels of observed data, ranging from 50% to 90%. Detailed experiment protocols and additional results are included in the supplementary material.
Classification of Irregularly-Sampled Data
We utilized the PhysioNet Sepsis dataset (Reyna et al. 2019), comprising 40,335 patient instances with 34 temporal variables. Following Kidger et al. (2020), we addressed data irregularity using two classification approaches: with and without observation intensity (OI). Observation intensity (OI) indicates patient condition severity, augmenting each time series point with an intensity index when utilized. Model performance was evaluated using Area Under the Receiver Operating Characteristic Curve (AUROC), with benchmark comparisons from Jhin et al. (2022).
Methods Test AUROC Memory (MB)
OI No OI OI No OI
GRU-$\Delta t$ 0.878 ± 0.006 0.840 ± 0.007 837 826
GRU-D 0.871 ± 0.022 0.850 ± 0.013 889 878
GRU-ODE 0.852 ± 0.010 0.771 ± 0.024 454 273
ODE-RNN 0.874 ± 0.016 0.833 ± 0.020 696 686
Latent-ODE 0.787 ± 0.011 0.495 ± 0.002 133 126
ACE-NODE 0.804 ± 0.010 0.514 ± 0.003 194 218
Neural CDE 0.880 ± 0.006 0.776 ± 0.009 244 122
ANCDE 0.900 ± 0.002 0.823 ± 0.003 285 129
EXIT 0.913 ± 0.002 0.836 ± 0.003 257 127
Proposed method 0.918 ± 0.003 0.873 ± 0.004 453 233
Table 4: AUROC on PhysioNet Sepsis dataset
In Table 4, our method demonstrates superior AUROC scores in both scenarios. Consistent with prior experimental findings, models based on flow exhibit enhanced efficacy relative to their non-flow neural network counterparts. These results confirm the effectiveness of our proposed methodologies in addressing challenges associated with time series irregularity and missingness.
Forecasting with the Partial Observations
We utilized two datasets, MuJoCo and Google, with different levels of observation for the forecasting task. Detailed explanations of each dataset, implementation details, and results are provided in the supplementary material.
MuJoCo.
We utilized the MuJoCo dataset (Todorov, Erez, and Tassa 2012), based on the Hopper configuration from the DeepMind control suite (Tassa et al. 2018). Following Jhin et al. (2023, 2024), we used 50 initial points to predict the next 10, introducing 30%, 50%, and 70% missing data scenarios. Table 5 illustrates the forecasting performance across diverse conditions, with our method consistently demonstrating minimal MSE in all settings.
Methods Test MSE
Regular 30% Dropped 50% Dropped 70% Dropped
GRU-$\Delta t$ 0.223 ± 0.020 0.198 ± 0.036 0.193 ± 0.015 0.196 ± 0.028
GRU-D 0.578 ± 0.042 0.608 ± 0.032 0.587 ± 0.039 0.579 ± 0.052
GRU-ODE 0.856 ± 0.016 0.857 ± 0.015 0.852 ± 0.015 0.861 ± 0.015
ODE-RNN 0.328 ± 0.225 0.274 ± 0.213 0.237 ± 0.110 0.267 ± 0.217
Latent-ODE 0.029 ± 0.011 0.056 ± 0.001 0.055 ± 0.004 0.058 ± 0.003
Augmented-ODE 0.055 ± 0.004 0.056 ± 0.004 0.057 ± 0.005 0.057 ± 0.005
ACE-NODE 0.039 ± 0.003 0.053 ± 0.007 0.053 ± 0.005 0.052 ± 0.006
Neural CDE 0.028 ± 0.002 0.027 ± 0.000 0.027 ± 0.001 0.026 ± 0.001
ANCDE 0.026 ± 0.001 0.025 ± 0.001 0.025 ± 0.001 0.024 ± 0.001
EXIT 0.026 ± 0.000 0.025 ± 0.004 0.026 ± 0.000 0.026 ± 0.001
LEAP 0.022 ± 0.002 0.022 ± 0.001 0.022 ± 0.002 0.022 ± 0.001
Proposed method 0.006 ± 0.001 0.008 ± 0.001 0.008 ± 0.001 0.008 ± 0.001
Table 5: Forecasting performance comparison on the MuJoCo dataset with varying percentages of data dropped
Figure 4: Forecasting performance versus the number of parameters on MuJoCo dataset with regular scenario (different flow model (with remarked *) versus conventional MLP)
Google.
We used the Google stock data (2011-2021)(Jhin et al. 2022), forecasting 5 price metrics for the next 10 days based on 50 days of history. Following Jhin et al. (2022)’s protocol, we introduced 30%, 50%, and 70% missing data scenarios. Table 6 presents the outcomes of forecasting stock data with partial observations. When comparing with benchmark methods, our method consistently presents lower MSE in the four different scenarios.
Methods Test MSE
Regular 30% Dropped 50% Dropped 70% Dropped
GRU-$\Delta t$ 0.0406 ± 0.002 0.0043 ± 0.003 0.0041 ± 0.002 0.0041 ± 0.001
GRU-D 0.0040 ± 0.004 0.0058 ± 0.018 0.0039 ± 0.001 0.0040 ± 0.001
GRU-ODE 0.0028 ± 0.016 0.0029 ± 0.001 0.0031 ± 0.004 0.0030 ± 0.002
ODE-RNN 0.0311 ± 0.044 0.0318 ± 0.066 0.0322 ± 0.056 0.0324 ± 0.053
Latent-ODE 0.0030 ± 0.011 0.0032 ± 0.001 0.0033 ± 0.002 0.0032 ± 0.003
Augmented-ODE 0.0023 ± 0.002 0.0025 ± 0.002 0.0029 ± 0.006 0.0023 ± 0.005
ACE-NODE 0.0022 ± 0.003 0.0024 ± 0.001 0.0022 ± 0.002 0.0025 ± 0.002
Neural CDE 0.0037 ± 0.062 0.0038 ± 0.073 0.0032 ± 0.035 0.0039 ± 0.038
ANCDE 0.0020 ± 0.002 0.0022 ± 0.001 0.0021 ± 0.002 0.0021 ± 0.001
EXIT 0.0016 ± 0.002 0.0016 ± 0.001 0.0017 ± 0.001 0.0019 ± 0.003
Proposed method 0.0010 ± 0.0001 0.0010 ± 0.0001 0.0009 ± 0.0001 0.0013 ± 0.0002
Table 6: Forecasting performance comparison on the Google dataset with varying percentages of data dropped
Figure 5: Forecasting performance versus the number of parameters on Google dataset with regular scenario (different flow model (with remarked *) versus conventional MLP)
Discussion
Performance versus Network Capacity.
Our analysis reveals that the performance improvements of DualDynamics are not merely a result of increased network capacity, but rather stem from the synergistic integration of implicit and explicit methods in our framework.
To validate DualDynamics’ effectiveness, we conducted controlled experiments comparing it against models with similar parameter counts. Figures 4 and 5 illustrate the relationship between parameter count and MSE for our flow architecture versus conventional MLPs. Figure 4 shows that increasing MLP capacity yields limited improvement on the MuJoCo dataset, while DualDynamics maintains superior performance. Figure 5 demonstrates that with optimal flow configuration, DualDynamics consistently outperforms MLPs on the Google dataset across all model sizes.
This performance advantage persists even when adjusting baseline model complexity to match DualDynamics, indicating that our framework’s superiority stems from its architectural design rather than network capacity. These results highlight DualDynamics’ efficiency in capturing complex temporal dynamics in irregular time series.
(a) Physionet Mortality
(b) Physionet Sepsis
(c) MuJoCo
(d) Google
Figure 6: Ablation study of different flow model (with remarked *) versus conventional MLP
Ablation Study of Flow Model Configurations.
Figure 6 presents a comprehensive overview of our ablation study with four datasets, illustrating the performance of various flow configurations across different tasks, as detailed in the preceding sections. The results demonstrate a clear advantage of flow-based architectures over the conventional MLP approach, refer to the detailed results in the supplementary material. These results suggest that incorporating implicit and explicit could lead to significant improvements in model performance and capabilities.
Variations in Our Framework.
Our framework’s architecture allows for several potential variations, each with distinct trade-offs. One possible modification is to utilize only ${\bm{z}}(t)$ instead of $\hat{{\bm{z}}}(t)$ after training the unified framework. However, this approach significantly limits the model’s expressiveness, as it fails to leverage the unique properties of the flow model. Another consideration is the replacement of the flow model with more sophisticated architectures, such as feedforward networks (Hornik, Stinchcombe, and White 1989), gated recurrent units (Chung et al. 2014), temporal convolutional networks (Lea et al. 2017), or transformer-based attention mechanisms (Vaswani et al. 2017). For further details and quantitative comparisons, we refer the reader to the supplementary material.
Conclusion
We introduce DualDynamics, an innovative methodology that integrates implicit and explicit methods for modeling irregularly-sampled time series data. This sophisticated paradigm excels in capturing intricate temporal characteristics inherent in continuous-time processes, offering an enhanced framework for representation learning in time series analysis. By synergistically combining these approaches, our method strikes a balance between expressive power and computational efficiency, addressing key challenges in temporal modeling for the irregular time series.
Our approach consistently surpasses existing benchmarks in classification, interpolation, and forecasting tasks, significantly elevating the precision and depth of understanding in temporal dynamics. These outstanding results highlight DualDynamics’ potential for diverse real-world applications. Its adaptability to various temporal modeling challenges positions it as a significant advancement in the irregular time series analysis in practices.
Ethics Statement
We commit to conducting our research with integrity, ensuring ethical practices and the responsible use of technology, fully aligning our efforts with established academic and scientific standards to promote trust and accountability.
Acknowledgments
We thank the teams and individuals for their efforts in the real-world dataset preparation and curation for our research, especially the UEA & UCR repository for the numerous datasets that we extensively analyzed.
This research was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2024-00407852); Korea Health Technology R&D Project through the Korea Health Industry Development Institute (KHIDI), funded by the Ministry of Health and Welfare, Republic of Korea (HI19C1095); the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.2020-0-01336, Artificial Intelligence Graduate School Program (UNIST)); and the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2021-II212068, Artificial Intelligence Innovation Hub).
Appendix A Experimental Details
We consider following experiments: (1) Robustness to Dataset Shift, (2) Classification of irregularly-sampled data, (3) Interpolation of Missing Data, and (4) Forecasting with the partial observations. All experiments were performed using a server on Ubuntu 22.04 LTS, equipped with an Intel(R) Xeon(R) Gold 6242 CPU and multiple NVIDIA A100 40GB GPUs. The source code for our experiments and datasets can be accessed at https://github.com/yongkyung-oh/DualDynamics.
Figure 7 illustrates the architecture of our proposed DualDynamics. This framework integrates implicit and explicit learning components for enhanced time series modeling. The implicit component can be implemented using Neural ODEs, Neural CDEs, or Neural SDEs, each adept at capturing complex temporal dependencies. The explicit component, designed to refine these representations, utilizes advanced neural flow models: ResNet Flow, GRU Flow, or Coupling Flow. This combination of implicit and explicit methods in DualDynamics aims to balance flexibility in modeling irregular time series with computational efficiency, potentially improving performance.
Figure 7: The detailed architecture of the proposed DualDynamics with Neural CDE
Implicit Component: NDE-Based Approaches
Neural ODEs
Let ${\bm{x}}=(x_{0},x_{1},\ldots,x_{n})$ be a vector of original (possibly, irregularly-sampled) observations. For a given time $t$, let us denote a latent state as ${\bm{z}}(t)$. Neural ODEs can be formally expressed as:
${\bm{z}}(t)$ $={\bm{z}}(0)+\int_{0}^{t}f(s,{\bm{z}}(s);\theta_{f}){\textnormal{d}}s,$ (20)
with the initial condition ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$ where $h:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{z}}$ is an affine function with parameter $\theta_{h}$. In Eq. (20), the function $h(\cdot;\theta_{h})$ can capture the intrinsic features from the input to the latent state. $f(s,{\bm{z}}(s);\theta_{f})$, which approximates the rate of change, $\dot{{\bm{z}}}\equiv\frac{{\textnormal{d}}{\bm{z}}(t)}{{\textnormal{d}}t}$, is modelled using a neural network parameterized by $\theta_{f}$. However, the performance of Neural ODEs often diminishes due to their dependence on initial conditions (Kidger 2022).
Neural CDEs
Neural CDEs provide a framework for constructing a continuous-time representation by formulating a control path $X(t)$, generated through natural cubic spline interpolation of observed data (Kidger et al. 2020). The Neural CDEs is expressed via the Riemann–Stieltjes integral:
${\bm{z}}(t)$ $={\bm{z}}(0)+\int_{0}^{t}f(s,{\bm{z}}(s);\theta_{f})\,{\textnormal{d}}X(s),$ (21)
with the initial condition ${\bm{z}}(0)=h(x_{0};\theta_{h})$ where $f(s,{\bm{z}}(s);\theta_{f})$ provides an approximation to $\frac{{\textnormal{d}}{\bm{z}}(t)}{{\textnormal{d}}X(t)}$, differentiating Neural CDEs from Neural ODEs in their approximation approach. To evaluate the integral in Eq. (21), one may utilize the conventional ODE solvers by recognizing that $\dot{{\bm{z}}}(t)\equiv\frac{{\textnormal{d}}{\bm{z}}(t)}{{\textnormal{d}}X(t)}$.
Neural SDEs
Neural SDEs extend the concept of Neural ODEs by incorporating a stochastic term to model uncertainty and noise in the dynamics (Oh, Lim, and Kim 2024). For a given time $t$ and latent state ${\bm{z}}(t)$, Neural SDEs can be formally expressed as:
${\bm{z}}(t)$ $={\bm{z}}(0)+\int_{0}^{t}f(s,{\bm{z}}(s);\theta_{f}){\textnormal{d}}s+\int_{0}^{t}g(s,{\bm{z}}(s);\theta_{g}){\textnormal{d}}W(s),$ (22)
with the initial condition ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$ where $h:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{z}}$ is an affine function with parameter $\theta_{h}$, similar to Neural ODEs. Here, $f(t,{\bm{z}}(t);\theta_{f})$ represents the drift term, which models the deterministic part of the dynamics, while $g(t,{\bm{z}}(t);\theta_{g})$ is the diffusion term that captures the stochastic component. $W(t)$ denotes a standard Wiener process (or Brownian motion). In Eq. (22), both $f$ and $g$ are modeled using neural networks parameterized by $\theta_{f}$ and $\theta_{g}$, respectively. Neural SDEs provide a framework for modeling complex, stochastic temporal dynamics, allowing for the representation of uncertainty in the latent state evolution as continues-time perspective.
Explicit Component: Flow Model Configurations
We integrate three distinct flow architectures proposed in Biloš et al. (2021) within the proposed framework.
ResNet Flow.
Originating from the concept of Residual Networks (ResNets) (He et al. 2016), the ResNet flow model can be conceptualized as a continuous-time analogue of ResNets. The core dynamics of ${\bm{z}}(t)$ in ResNet flow are described by the equation:
${\mathcal{G}}(t,{\bm{z}})={\bm{z}}+\varphi(t)g(t,{\bm{z}}),$ (23)
where $\varphi$ represents a temporal embedding function and $g$ denotes a continuous, nonlinear transformation parameterized by $\theta_{g}$. To address the inherent non-invertibility in standard ResNets, spectral normalization techniques (Gouk et al. 2021) are employed, ensuring a bounded Lipschitz constant and thus, invertibility (Behrmann et al. 2019).
GRU Flow.
The GRU flow model extends the Gated Recurrent Unit (GRU) architecture (Chung et al. 2014) to a continuous-time setting. In this model, the hidden state ${\bm{z}}$ evolves continuously according to an ordinary differential equation, akin to the updates in traditional discrete GRUs. De Brouwer et al. (2019) introduced GRU-ODE, and Biloš et al. (2021) further refined it to an invertible form within the Neural Flow context, as expressed by:
${\mathcal{G}}(t,{\bm{z}})={\bm{z}}+\varphi(t)(1-g_{1}(t,{\bm{z}}))\odot(g_{2}(t,{\bm{z}})-{\bm{z}}),$ (24)
where $\varphi(t)$ is a time-dependent embedding function, and $g_{1}$, $g_{2}$ are adapted from the GRU, where $g_{1}(t,{\bm{z}})=\alpha\cdot\texttt{sigmoid}(f_{1}(t,{\bm{z}}))$, and $g_{2}(t,{\bm{z}})=\texttt{tanh}(f_{2}(t,g_{3}(t,{\bm{z}})\odot{\bm{z}})$, and $g_{3}(t,{\bm{z}})=\beta\cdot\texttt{sigmoid}(f_{3}(t,{\bm{z}})$. $f_{1}$, $f_{2}$, $f_{3}$ are any arbitrary neural networks.
Coupling Flow.
The concept of Coupling flow, initially introduced by Dinh, Krueger, and Bengio (2014) and Dinh, Sohl-Dickstein, and Bengio (2016), involves a bijective transformation characterized by its computational efficiency and flexibility. Input dimensions are partitioned into two disjoint subsets $d_{1}$ and $d_{2}$, leading to the transformation:
${\mathcal{G}}(t)_{d_{1}}={\bm{z}}_{d_{1}}\exp(u(t,{\bm{z}}_{d_{2}})\varphi_{u}(t))+v(t,{\bm{z}}_{d_{2}})\varphi_{v}(t),$ (25)
with $u$ and $v$ as neural networks, and $\varphi_{u}$, $\varphi_{v}$ as time-dependent embedding functions. Two disjoint set $d_{1}$ and $d_{2}$ satisfies $d_{1}\bigcup d_{2}=\left\{1,2,\ldots,d_{z}\right\}$. This transformation is inherently invertible, maintaining the topological properties of the state space.
Each flow model exhibits its own strengths and limitations. The choice of a flow model should be contingent upon the dataset’s nature, the problem’s specific requirements, and computational constraints. Therefore, the selection of a flow model is empirically refined during hyperparameter tuning.
Training Procedure
We present the algorithm for the proposed DualDynamics framework. The presented algorithm 1 concisely captures the training procedure of the DualDynamics framework. Here we showcase the example of proposed method using Neural CDE for classification task with cross-entropy loss.
Algorithm 1 Training Procedure for DualDynamics (using Neural CDE for classification task)
1:Initialized model parameters $\Theta=\{\theta_{h},\theta_{f},\theta_{k},\theta_{\mathcal{G}},\theta_{MLP}\}$.
2:Forward Pass:
3:    Compute initial latent state ${\bm{z}}(0)=h({\bm{x}}_{0};\theta_{h})$ and solve Neural CDE to obtain ${\bm{z}}(t)$:
${\bm{z}}(t)={\bm{z}}(0)+\int_{0}^{t}f(s,{\bm{z}}(s);\theta_{f})\,{\textnormal{d}}X(s)$
4:    Aggregate ${\bm{z}}(t)$ to compute $\hat{{\bm{z}}}(0)=k({\bm{z}};\theta_{k})$ and compute $\hat{{\bm{z}}}(t)$ using flow model ${\mathcal{G}}$:
$\hat{{\bm{z}}}(t)={\mathcal{G}}(t,\hat{{\bm{z}}}(0);\theta_{\mathcal{G}})$
5:Loss Computation:
6:    Compute cross-entropy loss $\mathcal{L}$ using $\hat{\bm{z}}(T)$ and target $y$.
$\mathcal{L}=-\sum_{i=1}^{C}y_{i}\log(\text{softmax}(MLP(\hat{\bm{z}}(T);\theta_{MLP}))_{i})$
7:Backward Pass:
8:    Compute gradients $\nabla_{\theta}\mathcal{L}$ using adjoint-based backpropagation:
$\frac{\partial\mathcal{L}}{\partial\theta_{f}}=\int_{0}^{T}\bm{\lambda}_{{\bm{z}}}(t)^{\top}\frac{\partial f(t,{\bm{z}}(t);\theta_{f})}{\partial\theta_{f}}\,\frac{{\textnormal{d}}X(t)}{{\textnormal{d}}t}{\textnormal{d}}t,\qquad\frac{\partial\mathcal{L}}{\partial\theta_{\mathcal{G}}}=\bm{\lambda}_{\hat{{\bm{z}}}}(T)^{\top}\frac{\partial{\mathcal{G}}(T,\hat{{\bm{z}}}(0);\theta_{{\mathcal{G}}})}{\partial\theta_{\mathcal{G}}}.$
9:    Update all parameters. $\Theta\leftarrow\Theta-\eta\nabla_{\Theta}\mathcal{L}$
Appendix B Detailed Results for “Robustness to Dataset Shift”
Data Preparation
We conducted classification experiments across 18 diverse datasets, in three distinct domains, from the University of East Anglia (UEA) and the University of California Riverside (UCR) Time Series Classification Repository  [^3] http://www.timeseriesclassification.com/ (Bagnall et al. 2018; Dau et al. 2019) using the python library sktime (Löning et al. 2019), as outlined in Table 7. Addressing the challenge posed by varying lengths of time series, the strategy of uniform scaling was employed to align all series to the dimension of the longest series (Keogh 2003; Yankov et al. 2007; Gao and Lin 2018; Tan et al. 2019). The partitioning of the dataset into training, validation, and testing subsets was executed in proportions of 0.70, 0.15, and 0.15 respectively. Subsequent to this partitioning, a random assortment of missing observations was introduced for each variable. Lastly, these modified variables are integrated into combined dataset.
Domain Dataset
Total number of
samples
Number of
classes
Dimension of
time series
Length of
time series
Motion & HAR BasicMotions 80 4 6 100
Epilepsy 275 4 3 206
PickupGestureWiimoteZ 100 10 1 29-361
ShakeGestureWiimoteZ 100 10 1 41-385
ToeSegmentation1 268 2 1 277
ToeSegmentation2 166 2 1 343
ECG & EEG Blink 950 2 4 510
ECG200 200 2 1 96
SelfRegulationSCP1 561 2 6 896
SelfRegulationSCP2 380 2 7 1152
StandWalkJump 27 3 4 2500
TwoLeadECG 1162 2 1 82
Sensor DodgerLoopDay 158 7 1 288
DodgerLoopGame 158 2 1 288
DodgerLoopWeekend 158 2 1 288
Lightning2 121 2 1 637
Lightning7 143 7 1 319
Trace 200 4 1 275
Table 7: Description of datasets for the classification task
Experimental Protocol
We utilized the source codes for torchcde library [^4] https://github.com/patrick-kidger/torchcde (Kidger et al. 2020; Morrill et al. 2021),and Neural Flow [^5] https://github.com/mbilos/neural-flows-experiments (Biloš et al. 2021) for the proposed method. In case of benchmark, we used implementation of RNN, LSTM, and GRU using pytorch [^6] https://pytorch.org/ (Paszke et al. 2019). Also, this research incorporated the primary Neural CDE algorithms from their original repository [^7] https://github.com/patrick-kidger/NeuralCDE (Kidger et al. 2020), which included methodologies such as GRU-$\Delta t$, GRU-D, GRU-ODE, ODE-RNN, and Neural CDE. Further, this study leveraged the primary source code of MTAN [^8] https://github.com/reml-lab/mTAN (Shukla and Marlin 2021), MIAM [^9] https://github.com/ku-milab/MIAM (Lee et al. 2022), ODE-LSTM [^10] https://github.com/mlech26l/ode-lstms (Lechner and Hasani 2020), Neural RDE [^11] https://github.com/jambo6/neuralRDEs (Morrill et al. 2021), ANCDE [^12] https://github.com/sheoyon-jhin/ANCDE (Jhin et al. 2024), EXIT [^13] https://github.com/sheoyon-jhin/EXIT (Jhin et al. 2022), and LEAP [^14] https://github.com/alflsowl12/LEAP (Jhin et al. 2023) for a comprehensive evaluation.
The integrity of the comparative analysis was upheld by utilizing the original architectural designs for all methods. Due to the variability of optimal hyperparameters across different methods and datasets, the ray library [^15] https://github.com/ray-project/ray (Moritz et al. 2018; Liaw et al. 2018) was employed for the refinement of these parameters. This library facilitates the automatic tuning of hyperparameters to minimize validation loss, marking a departure from previous studies that necessitated manual adjustments per dataset and model. Optimal hyperparameters determined for regular time series datasets were also applied to their irregular counterparts. The Euler(-Maruyama) method was consistently used as the ODE solver for all approaches in implicit methods.
In each model and dataset, the hyperparameters fine-tuned to minimize validation loss included the number of layers $n_{l}$ and the dimensions of the hidden vector $n_{h}$. For RNN-based methods such as RNN, LSTM, and GRU, these hyperparameters were integral to the fully-connected layer, while for NDE-based approaches, they were pertinent to the embedding layer and vector fields. Hyperparameter optimization was systematically conducted: the learning rate $lr$ was calibrated from $10^{-4}$ to $10^{-1}$ via log-uniform search; $n_{l}$ was set through grid search from the set $\left\{1,2,3,4\right\}$; and $n_{h}$ was determined from the set $\left\{16,32,64,128\right\}$ through grid search. The batch size was chosen from the set $\left\{16,32,64,128\right\}$, considering the total size of the data. All models underwent a training duration of 100 epochs, with the most effective model being chosen based on the lowest validation loss, thereby ensuring its generalizability.
Motivating Example: BasicMotions
Table 8 presents the classification outcomes for both with and without missingness. Computation times are measured with regular setting using a single GPU. While Neural CDE generally exhibits robust performance, its extended variants tend to require increased computation time and do not consistently perform well across a broad range of tasks compared to the benchmark performance. On the other hand, Neural Flow achieves rapid convergence with reduced performance. In contrast, our proposed method not only shows superior performance but also maintains a balanced computational demand.
Methods Regular 30% Missing 50% Missing 70% Missing Average Computation time (s)
RNN -based RNN 0.750 ± 0.059 0.633 ± 0.095 0.600 ± 0.070 0.583 ± 0.059 0.642 ± 0.071 2.27 ± 0.06
LSTM 0.833 ± 0.132 0.833 ± 0.156 0.700 ± 0.095 0.700 ± 0.112 0.767 ± 0.124 2.84 ± 0.93
GRU 0.950 ± 0.046 0.967 ± 0.075 0.900 ± 0.091 0.883 ± 0.173 0.925 ± 0.096 2.97 ± 0.05
GRU-$\Delta t$ 0.950 ± 0.075 1.000 ± 0.000 0.967 ± 0.046 0.950 ± 0.046 0.967 ± 0.041 34.39 ± 0.27
GRU-D 0.967 ± 0.075 0.950 ± 0.112 0.933 ± 0.070 0.967 ± 0.046 0.954 ± 0.075 42.43 ± 0.29
Attention -based Transformer 0.983 ± 0.037 0.900 ± 0.109 0.967 ± 0.046 0.933 ± 0.091 0.946 ± 0.071 5.75 ± 0.39
MTAN 0.950 ± 0.112 0.717 ± 0.247 0.717 ± 0.247 0.700 ± 0.046 0.771 ± 0.163 38.77 ± 0.71
MIAM 0.967 ± 0.046 0.983 ± 0.037 0.933 ± 0.109 0.867 ± 0.126 0.938 ± 0.079 112.34 ± 1.43
ODE -based GRU-ODE 1.000 ± 0.000 0.967 ± 0.075 0.983 ± 0.037 0.933 ± 0.070 0.971 ± 0.045 519.78 ± 0.34
ODE-RNN 0.983 ± 0.037 1.000 ± 0.000 1.000 ± 0.000 1.000 ± 0.000 0.996 ± 0.009 113.29 ± 0.17
ODE-LSTM 0.733 ± 0.109 0.683 ± 0.037 0.717 ± 0.126 0.567 ± 0.216 0.675 ± 0.122 73.03 ± 0.73
CDE -based Neural CDE 0.983 ± 0.037 1.000 ± 0.000 0.983 ± 0.037 0.900 ± 0.037 0.967 ± 0.028 112.53 ± 0.41
Neural RDE 0.983 ± 0.037 0.983 ± 0.037 0.950 ± 0.075 0.833 ± 0.083 0.938 ± 0.058 93.79 ± 0.44
ANCDE 1.000 ± 0.000 0.967 ± 0.046 0.983 ± 0.037 0.900 ± 0.137 0.963 ± 0.055 360.50 ± 0.47
EXIT 0.300 ± 0.112 0.450 ± 0.095 0.417 ± 0.102 0.367 ± 0.139 0.383 ± 0.112 334.53 ± 0.28
LEAP 0.317 ± 0.149 0.300 ± 0.095 0.317 ± 0.124 0.167 ± 0.083 0.275 ± 0.113 163.04 ± 0.59
Flow-based Neural Flow 0.717 ± 0.139 0.433 ± 0.207 0.433 ± 0.109 0.367 ± 0.126 0.487 ± 0.145 4.64 ± 0.28
Proposed method 1.000 ± 0.000 1.000 ± 0.000 0.983 ± 0.037 0.983 ± 0.037 0.992 ± 0.019 119.08 ± 2.70
Table 8: Performance and computation time comparison using ‘BasicMotions’ dataset under regular (0% missingness) and irregular (30%, 50%, and 70% missingness) scenarios. Computation time is calculated without early-stopping.
Variations in model architecture
Our framework’s architecture allows for several potential variations, each with distinct trade-offs. In this section, we compare the versatility of our framework with two potential perspectives. One possible modification is to utilize only ${\bm{z}}(t)$ instead of $\hat{{\bm{z}}}(t)$ after training the unified framework. However, this approach would significantly limit the model’s expressiveness, as it fails to leverage the unique properties of the flow model, particularly its ability to model complex probability distributions through invertible transformations, as shown in Table 9.
Another consideration is the replacement of the flow model with more sophisticated architectures such as feedforward networks (Hornik, Stinchcombe, and White 1989), gated recurrent unit (Chung et al. 2014), temporal convolution networks (Lea et al. 2017), or transformer-based attention mechanisms (Vaswani et al. 2017). Table 10 summarizes the analysis of different model components in our framework. While these alternatives demonstrate strong performance on regularly-sampled time series, they often struggle with irregular sampling patterns and missing data. Compared to that, the flow model in our current architecture specifically addresses these challenges. The flow model in our current architecture specifically addresses these challenges through its continuous and invertible nature, making it particularly well-suited for handling irregular time series data with missing observations.
Configuration Regular 30% Missing 50% Missing 70% Missing
Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Proposed method
(Neural CDE)
1.000 (0.000) 0.007 (0.008) 1.000 (0.000) 0.021 (0.027) 0.983 (0.037) 0.044 (0.063) 0.983 (0.037) 0.069 (0.141)
– using ${\bm{z}}(t)$ instead 1.000 (0.000) 0.152 (0.087) 1.000 (0.000) 0.144 (0.119) 1.000 (0.000) 0.126 (0.126) 0.883 (0.095) 0.346 (0.174)
Neural CDE 0.983 (0.037) 0.037 (0.056) 1.000 (0.000) 0.048 (0.043) 0.983 (0.037) 0.071 (0.061) 0.900 (0.037) 0.433 (0.242)
Neural Flow 0.717 (0.139) 0.969 (0.340) 0.433 (0.207) 1.384 (0.285) 0.433 (0.109) 1.382 (0.382) 0.367 (0.126) 1.521 (0.343)
Table 9: Comparative study of latent value selection using ‘BasicMotions’ dataset
Configuration Regular 30% Missing 50% Missing 70% Missing
Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Proposed method
(Neural CDE)
1.000 (0.000) 0.007 (0.008) 1.000 (0.000) 0.021 (0.027) 0.983 (0.037) 0.044 (0.063) 0.983 (0.037) 0.069 (0.141)
[…]
– Conventional MLP 0.950 (0.075) 0.131 (0.104) 0.933 (0.037) 0.263 (0.179) 0.950 (0.046) 0.167 (0.113) 0.933 (0.037) 0.214 (0.053)
– ResNet Flow 0.983 (0.037) 0.048 (0.053) 0.967 (0.075) 0.078 (0.121) 0.917 (0.102) 0.135 (0.121) 1.000 (0.000) 0.041 (0.035)
– GRU Flow 0.967 (0.046) 0.093 (0.073) 0.950 (0.075) 0.116 (0.107) 0.967 (0.046) 0.158 (0.096) 0.983 (0.037) 0.088 (0.057)
– Coupling Flow 1.000 (0.000) 0.007 (0.008) 1.000 (0.000) 0.021 (0.027) 0.983 (0.037) 0.044 (0.063) 0.983 (0.037) 0.069 (0.141)
– Feedforward network 1.000 (0.000) 0.023 (0.014) 0.983 (0.037) 0.066 (0.042) 0.900 (0.109) 0.314 (0.278) 0.850 (0.070) 0.432 (0.152)
– Gated recurrent unit 0.983 (0.037) 0.050 (0.063) 0.883 (0.075) 0.402 (0.339) 0.883 (0.046) 0.333 (0.176) 0.867 (0.075) 0.450 (0.199)
– Temporal convolution 0.850 (0.149) 0.386 (0.298) 0.850 (0.091) 0.455 (0.383) 0.900 (0.070) 0.408 (0.211) 0.767 (0.109) 0.575 (0.272)
– Transformer 0.917 (0.083) 0.254 (0.252) 0.950 (0.112) 0.292 (0.639) 0.950 (0.075) 0.195 (0.286) 0.867 (0.095) 0.521 (0.390)
Table 10: Comparative study of component selection using ‘BasicMotions’ dataset
Configuration Regular 30% Missing 50% Missing 70% Missing
Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Proposed method
(Neural CDE)
0.724 (0.090) 0.631 (0.145) 0.720 (0.088) 0.703 (0.180) 0.691 (0.091) 0.712 (0.173) 0.697 (0.098) 0.698 (0.176)
[…]
– Conventional MLP 0.592 (0.084) 0.812 (0.143) 0.578 (0.091) 0.870 (0.152) 0.578 (0.087) 0.888 (0.163) 0.582 (0.085) 0.879 (0.136)
– ResNet Flow 0.678 (0.084) 0.714 (0.141) 0.662 (0.094) 0.791 (0.181) 0.638 (0.086) 0.791 (0.160) 0.625 (0.100) 0.812 (0.138)
– GRU Flow 0.669 (0.079) 0.685 (0.135) 0.662 (0.078) 0.776 (0.184) 0.656 (0.074) 0.803 (0.178) 0.652 (0.083) 0.810 (0.191)
– Coupling Flow 0.675 (0.098) 0.715 (0.166) 0.673 (0.090) 0.745 (0.197) 0.667 (0.096) 0.769 (0.185) 0.662 (0.102) 0.758 (0.194)
Proposed method
(Neural ODE)
0.540 (0.059) 0.948 (0.064) 0.534 (0.076) 0.971 (0.078) 0.535 (0.064) 0.977 (0.077) 0.531 (0.066) 0.978 (0.070)
[…]
– Conventional MLP 0.521 (0.062) 0.957 (0.065) 0.503 (0.067) 0.981 (0.082) 0.494 (0.063) 0.993 (0.085) 0.496 (0.071) 0.988 (0.071)
– ResNet Flow 0.528 (0.071) 0.960 (0.062) 0.505 (0.059) 0.978 (0.087) 0.503 (0.051) 0.987 (0.081) 0.504 (0.059) 0.987 (0.076)
– GRU Flow 0.525 (0.057) 0.965 (0.063) 0.530 (0.083) 0.974 (0.083) 0.520 (0.075) 0.999 (0.104) 0.515 (0.068) 0.986 (0.068)
– Coupling Flow 0.532 (0.061) 0.944 (0.064) 0.528 (0.077) 0.976 (0.075) 0.522 (0.057) 0.983 (0.072) 0.509 (0.062) 0.980 (0.068)
Proposed method
(Neural SDE)
0.543 (0.066) 0.958 (0.083) 0.531 (0.063) 0.971 (0.063) 0.537 (0.073) 0.974 (0.071) 0.541 (0.061) 0.970 (0.068)
[…]
– Conventional MLP 0.516 (0.065) 0.961 (0.084) 0.495 (0.070) 0.995 (0.075) 0.511 (0.073) 0.984 (0.074) 0.503 (0.068) 0.983 (0.073)
– ResNet Flow 0.526 (0.066) 0.965 (0.064) 0.518 (0.070) 0.964 (0.075) 0.513 (0.061) 0.972 (0.079) 0.522 (0.060) 0.973 (0.073)
– GRU Flow 0.520 (0.058) 0.980 (0.086) 0.528 (0.069) 0.980 (0.075) 0.522 (0.078) 0.982 (0.063) 0.515 (0.066) 0.984 (0.070)
– Coupling Flow 0.528 (0.071) 0.961 (0.087) 0.513 (0.068) 0.976 (0.076) 0.511 (0.062) 0.983 (0.083) 0.526 (0.071) 0.981 (0.070)
Table 11: Ablation study of the proposed method on 18 datasets under regular and three missing rates (Values in parentheses show average of 18 standard deviations)
Detailed Results of Ablation Study
Table 11 presents a comprehensive ablation study of our DualDynamics framework, comparing various combinations of implicit components (Neural ODE, CDE, and SDE) and explicit components (ResNet Flow, GRU Flow, and Coupling Flow) against a non-invertible conventional MLP baseline. Neural CDEs consistently outperform Neural ODEs and SDEs, aligning with previous findings in the literature. While the optimal flow configuration varies across datasets, all flow models demonstrate superior performance compared to the MLP baseline. This consistent improvement underscores the efficacy of our framework’s synergistic integration of implicit and explicit components.
Detailed Results Three Domains
Our investigation involved 18 datasets classified into three domains. We collate the outcomes across three domains: Motion & HAR, ECG & EEG, and Sensor datasets, in Tables 13, 13, and 14. We observe variability in the performance metrics due to the unique characteristics of each domain. Overall, the proposed methodology demonstrates superior performance relative to benchmark models, in the different domains and missing rate scenarios.
Methods Regular 30% Missing 50% Missing 70% Missing Average
Accuracy Rank Accuracy Rank Accuracy Rank Accuracy Rank Accuracy Rank
RNN 0.518 (0.062) 13.3 0.445 (0.077) 15.8 0.427 (0.061) 15.6 0.415 (0.070) 14.5 0.451 (0.067) 14.8
LSTM 0.568 (0.089) 11.6 0.522 (0.092) 11.9 0.494 (0.094) 12.3 0.467 (0.090) 13.3 0.513 (0.091) 12.3
GRU 0.696 (0.095) 7.2 0.629 (0.076) 10.3 0.612 (0.089) 9.6 0.631 (0.111) 7.8 0.642 (0.093) 8.7
GRU-$\Delta t$ 0.613 (0.076) 9.3 0.624 (0.079) 7.6 0.663 (0.064) 7.3 0.624 (0.077) 9.3 0.631 (0.074) 8.4
GRU-D 0.587 (0.077) 11.6 0.563 (0.089) 11.3 0.540 (0.056) 13.2 0.564 (0.054) 11.2 0.563 (0.069) 11.8
Transformer 0.772 (0.060) 3.2 0.703 (0.070) 4.8 0.718 (0.079) 4.3 0.689 (0.069) 4.3 0.720 (0.070) 4.1
MTAN 0.603 (0.096) 12.7 0.583 (0.116) 9.9 0.587 (0.099) 9.2 0.593 (0.063) 9.2 0.591 (0.093) 10.2
MIAM 0.590 (0.091) 9.0 0.603 (0.063) 7.8 0.589 (0.066) 8.6 0.562 (0.044) 9.8 0.586 (0.066) 8.8
GRU-ODE 0.662 (0.081) 8.8 0.647 (0.091) 9.5 0.656 (0.072) 8.7 0.657 (0.084) 7.0 0.656 (0.082) 8.5
ODE-RNN 0.621 (0.087) 9.8 0.603 (0.078) 8.6 0.588 (0.095) 8.8 0.627 (0.064) 7.6 0.610 (0.081) 8.7
ODE-LSTM 0.550 (0.083) 12.8 0.507 (0.080) 12.8 0.475 (0.102) 13.7 0.426 (0.094) 13.2 0.489 (0.089) 13.1
Neural CDE 0.751 (0.096) 5.9 0.728 (0.080) 5.5 0.717 (0.100) 5.9 0.738 (0.111) 6.2 0.733 (0.097) 5.9
Neural RDE 0.702 (0.091) 6.2 0.703 (0.077) 6.1 0.704 (0.088) 6.5 0.665 (0.095) 7.6 0.693 (0.088) 6.6
ANCDE 0.710 (0.068) 5.8 0.716 (0.075) 5.0 0.724 (0.074) 4.1 0.703 (0.077) 5.5 0.713 (0.074) 5.1
EXIT 0.548 (0.086) 10.8 0.584 (0.073) 9.3 0.567 (0.086) 10.4 0.534 (0.097) 11.8 0.558 (0.086) 10.6
LEAP 0.397 (0.084) 15.7 0.374 (0.086) 16.0 0.381 (0.077) 14.1 0.335 (0.062) 14.8 0.372 (0.077) 15.1
Neural Flow 0.527 (0.068) 11.3 0.420 (0.081) 13.3 0.389 (0.084) 14.1 0.360 (0.065) 14.3 0.424 (0.074) 13.2
Proposed method 0.704 (0.082) 6.3 0.702 (0.074) 5.6 0.696 (0.088) 4.9 0.723 (0.123) 3.7 0.706 (0.092) 5.1
Table 12: Average classification performance on 6 datasets in Motion & HAR
Methods Regular 30% Missing 50% Missing 70% Missing Average
Accuracy Rank Accuracy Rank Accuracy Rank Accuracy Rank Accuracy Rank
RNN 0.604 (0.075) 10.1 0.532 (0.058) 14.2 0.545 (0.062) 11.8 0.506 (0.056) 14.2 0.547 (0.063) 12.6
LSTM 0.609 (0.054) 10.8 0.592 (0.066) 8.5 0.568 (0.057) 9.7 0.546 (0.048) 10.1 0.579 (0.056) 9.8
GRU 0.692 (0.073) 5.8 0.704 (0.059) 5.2 0.658 (0.068) 7.9 0.661 (0.068) 6.9 0.679 (0.067) 6.5
GRU-$\Delta t$ 0.602 (0.047) 12.6 0.616 (0.052) 10.7 0.624 (0.062) 8.5 0.616 (0.069) 9.8 0.614 (0.057) 10.4
GRU-D 0.621 (0.056) 10.0 0.606 (0.066) 11.0 0.618 (0.055) 9.8 0.585 (0.061) 12.3 0.608 (0.060) 10.8
Transformer 0.693 (0.067) 8.9 0.643 (0.056) 10.2 0.621 (0.066) 10.8 0.621 (0.088) 11.1 0.645 (0.069) 10.3
MTAN 0.688 (0.081) 8.2 0.680 (0.115) 5.8 0.676 (0.093) 4.9 0.699 (0.093) 4.8 0.686 (0.096) 5.9
MIAM 0.634 (0.062) 11.3 0.651 (0.063) 7.9 0.644 (0.055) 7.2 0.637 (0.058) 7.8 0.642 (0.060) 8.5
GRU-ODE 0.650 (0.057) 8.1 0.661 (0.057) 6.8 0.653 (0.063) 7.5 0.633 (0.080) 8.4 0.649 (0.064) 7.7
ODE-RNN 0.664 (0.068) 6.3 0.650 (0.050) 7.4 0.644 (0.048) 9.0 0.647 (0.050) 6.2 0.651 (0.054) 7.2
ODE-LSTM 0.580 (0.058) 9.6 0.546 (0.065) 12.3 0.550 (0.054) 12.8 0.545 (0.036) 12.8 0.555 (0.053) 11.8
Neural CDE 0.643 (0.040) 8.1 0.632 (0.038) 8.3 0.625 (0.041) 8.8 0.596 (0.069) 8.3 0.624 (0.047) 8.3
Neural RDE 0.591 (0.065) 12.5 0.621 (0.060) 8.6 0.602 (0.070) 9.3 0.577 (0.073) 9.3 0.598 (0.067) 9.9
ANCDE 0.617 (0.075) 9.3 0.612 (0.060) 9.3 0.596 (0.059) 11.3 0.590 (0.072) 7.9 0.604 (0.066) 9.4
EXIT 0.599 (0.092) 10.8 0.552 (0.084) 12.8 0.563 (0.092) 11.2 0.574 (0.044) 10.2 0.572 (0.078) 11.2
LEAP 0.572 (0.051) 12.3 0.527 (0.080) 14.3 0.546 (0.093) 11.8 0.540 (0.086) 12.5 0.546 (0.077) 12.7
Neural Flow 0.565 (0.047) 11.7 0.547 (0.053) 12.4 0.522 (0.046) 13.3 0.516 (0.047) 12.4 0.538 (0.048) 12.5
Proposed method 0.708 (0.119) 4.8 0.747 (0.086) 5.5 0.703 (0.086) 5.5 0.683 (0.096) 6.3 0.710 (0.097) 5.5
Table 13: Average classification performance on 6 datasets in ECG & EEG
Methods Regular 30% Missing 50% Missing 70% Missing Average
Accuracy Rank Accuracy Rank Accuracy Rank Accuracy Rank Accuracy Rank
RNN 0.557 (0.078) 12.4 0.476 (0.089) 14.1 0.441 (0.123) 14.7 0.437 (0.079) 14.6 0.478 (0.092) 13.9
LSTM 0.588 (0.059) 10.5 0.542 (0.067) 10.2 0.484 (0.069) 12.3 0.503 (0.065) 10.4 0.529 (0.065) 10.8
GRU 0.633 (0.073) 9.0 0.585 (0.059) 9.7 0.563 (0.072) 9.5 0.526 (0.084) 11.0 0.577 (0.072) 9.8
GRU-$\Delta t$ 0.673 (0.072) 7.9 0.666 (0.075) 6.3 0.668 (0.080) 5.8 0.708 (0.075) 4.4 0.679 (0.076) 6.1
GRU-D 0.569 (0.130) 11.1 0.569 (0.105) 9.5 0.583 (0.115) 8.8 0.647 (0.070) 7.0 0.592 (0.105) 9.1
Transformer 0.695 (0.063) 8.3 0.643 (0.073) 9.2 0.669 (0.081) 7.1 0.618 (0.079) 9.3 0.656 (0.074) 8.4
MTAN 0.671 (0.088) 7.3 0.639 (0.063) 6.4 0.629 (0.074) 6.3 0.635 (0.063) 6.6 0.643 (0.072) 6.6
MIAM 0.493 (0.085) 15.3 0.500 (0.132) 14.2 0.482 (0.093) 14.1 0.432 (0.086) 15.5 0.477 (0.099) 14.8
GRU-ODE 0.677 (0.079) 6.0 0.674 (0.059) 5.8 0.684 (0.072) 4.3 0.687 (0.079) 5.2 0.680 (0.072) 5.3
ODE-RNN 0.673 (0.098) 5.9 0.642 (0.101) 7.3 0.647 (0.115) 5.7 0.684 (0.062) 5.8 0.662 (0.094) 6.2
ODE-LSTM 0.567 (0.083) 11.8 0.502 (0.061) 13.0 0.479 (0.049) 13.5 0.453 (0.075) 14.3 0.500 (0.067) 13.1
Neural CDE 0.650 (0.082) 8.8 0.657 (0.087) 8.4 0.641 (0.069) 7.3 0.623 (0.094) 7.3 0.643 (0.083) 8.0
Neural RDE 0.653 (0.089) 7.0 0.620 (0.076) 7.7 0.594 (0.077) 8.7 0.580 (0.069) 8.7 0.612 (0.078) 8.0
ANCDE 0.660 (0.106) 8.8 0.656 (0.115) 6.8 0.598 (0.108) 9.3 0.601 (0.069) 8.6 0.629 (0.100) 8.4
EXIT 0.637 (0.082) 8.7 0.605 (0.107) 9.5 0.606 (0.081) 8.6 0.584 (0.074) 9.2 0.608 (0.086) 9.0
LEAP 0.500 (0.052) 15.3 0.475 (0.044) 14.8 0.471 (0.054) 14.9 0.478 (0.073) 13.8 0.481 (0.056) 14.7
Neural Flow 0.542 (0.063) 12.8 0.487 (0.067) 14.5 0.454 (0.043) 15.5 0.437 (0.049) 15.5 0.480 (0.055) 14.6
Proposed method 0.760 (0.070) 4.3 0.712 (0.106) 3.8 0.673 (0.100) 4.9 0.687 (0.076) 3.9 0.708 (0.088) 4.2
Table 14: Average classification performance on 6 datasets in Sensor
Appendix C Detailed Results for “Interpolation of Missing Data”
The interpolation experiment was conducted using the 2012 PhysioNet Mortality dataset (Silva et al. 2012), which comprises multivariate time series data from ICU records. This dataset includes 37 variables collected from Intensive Care Unit (ICU) records, with each data instance capturing sporadic and infrequent measurements obtained within the initial 48-hour window post-admission to the ICU. Following the methodologies proposed by Rubanova, Chen, and Duvenaud (2019), we adjusted the observation timestamps to the nearest minute, thereby creating up to 2880 possible measurement intervals per time series. Our experiments adhered to the methodology proposed by Shukla and Marlin (2021), involving the alteration of observation frequency from 50% to 90% to predict the remaining data.
We followed experimental protocol suggested by Shukla and Marlin (2021), and its Github repository [^16] https://github.com/reml-lab/mTAN. For the proposed methodology, the training process spans $300$ epochs, employing a batch size of $64$ and a learning rate of $0.001$. To train our models on a dataset comprising irregularly sampled time series, we adopt a strategy from Shukla and Marlin (2021). This involves the modified VAE training method, where we optimize a normalized variational lower bound of the log marginal likelihood, grounded on the evidence lower bound (ELBO).
Configuration Test MSE ×$\times 10^{-3}$
50% 60% 70% 80% 90%
Conventional MLP 3.838 ± 0.021 3.655 ± 0.016 3.508 ± 0.010 3.316 ± 0.062 3.442 ± 0.078
ResNet Flow 3.655 ± 0.031 3.537 ± 0.020 3.373 ± 0.042 3.351 ± 0.154 3.171 ± 0.057
GRU Flow 3.754 ± 0.014 3.562 ± 0.019 3.470 ± 0.031 3.232 ± 0.101 3.222 ± 0.030
Coupling Flow 3.631 ± 0.049 3.659 ± 0.028 3.463 ± 0.032 3.224 ± 0.044 3.114 ± 0.050
Table 15: Ablation study on the PhysioNet Mortality Dataset MSE ×$\times 10^{-3}$)
Table 15 presents an ablation study on the interpolation task. Our DualDynamics framework significantly improves interpolation performance, with flow model architectures consistently outperforming conventional MLPs in terms of MSE. This performance gap, observed even with comparable network capacities, underscores the importance of our framework’s architectural design over mere network size in enhancing interpolation accuracy for time series data.
Hyperparameter optimization is conducted through a grid search, focusing on the number of layers $n_{l}\in\left\{1,2,3,4\right\}$ and hidden vector dimensions $n_{h}\in\left\{16,32,64,128\right\}$. The optimal hyperparameters for the proposed methods are remarked in bold within Table 16, encompassing observed data with 50%. We used the same hyperparameters to the remainder settings of 60%, 70%, 80%, and 90%.
$n_{l}$ $n_{h}$ Conventional MLP ResNet Flow GRU Flow Coupling Flow
1 16 4.574 ± 0.072 4.534 ± 0.041 4.753 ± 0.068 4.473 ± 0.159
32 4.332 ± 0.121 4.134 ± 0.014 4.241 ± 0.058 4.302 ± 0.074
64 3.991 ± 0.053 4.370 ± 0.062 3.901 ± 0.075 4.276 ± 0.145
128 3.839 ± 0.064 4.072 ± 0.042 4.097 ± 0.053 4.118 ± 0.083
2 16 4.593 ± 0.102 4.692 ± 0.032 4.685 ± 0.102 4.560 ± 0.056
32 4.253 ± 0.046 4.208 ± 0.104 4.540 ± 0.053 4.519 ± 0.106
64 4.459 ± 0.069 4.006 ± 0.062 4.104 ± 0.109 3.983 ± 0.047
128 3.876 ± 0.026 4.355 ± 0.043 3.778 ± 0.045 3.871 ± 0.037
3 16 4.580 ± 0.088 4.475 ± 0.116 4.866 ± 0.054 4.635 ± 0.129
32 4.332 ± 0.010 4.096 ± 0.111 4.253 ± 0.028 4.146 ± 0.030
64 3.997 ± 0.115 4.035 ± 0.153 4.185 ± 0.106 4.048 ± 0.119
128 3.838 ± 0.021 3.655 ± 0.031 3.754 ± 0.014 3.631 ± 0.049
4 16 4.552 ± 0.070 4.472 ± 0.077 4.744 ± 0.021 4.554 ± 0.067
32 4.312 ± 0.044 4.237 ± 0.054 4.696 ± 0.185 4.221 ± 0.065
64 4.082 ± 0.036 3.995 ± 0.023 4.175 ± 0.052 3.955 ± 0.054
128 3.848 ± 0.048 3.950 ± 0.042 3.982 ± 0.117 4.328 ± 0.048
Table 16: Results of hyperparameter tuning for PhysioNet Mortality Dataset (MSE ×$\times 10^{-3}$)
Appendix D Detailed Results for “Classification of Irregularly-Sampled Data”
We utilized the PhysioNet Sepsis dataset (Reyna et al. 2019) to investigate classification performance with irregularly-sampled time series. The dataset comprises 40,335 patient instances from ICU, characterized by 34 temporal variables such as heart rate, oxygen saturation, and body temperature. The objective is to detect the presence or absence of sepsis in each patient.
Given that the PhysioNet dataset represents irregular time series, with only 10% of the data points timestamped for each patient, the study addresses this irregularity through two distinct approaches to time series classification as suggested by Kidger et al. (2020): (i) classification utilizing observation intensity (OI) and (ii) classification without utilizing observation intensity (No OI). Observation intensity serves as an indicator of the severity of a patient’s condition, and when employed, each data point in the time series is supplemented with an index reflecting this intensity. Considering the dataset’s skewed distribution, performance is evaluated using the Area Under the Receiver Operating Characteristic Curve (AUROC) metric. Benchmark performance is established in Jhin et al. (2022).
Configuration Test AUROC Memory (MB)
OI No OI OI No OI
Conventional MLP 0.912 ± 0.011 0.861 ± 0.016 284 135
ResNet Flow 0.915 ± 0.003 0.875 ± 0.006 230 156
GRU Flow 0.916 ± 0.003 0.875 ± 0.004 355 178
Coupling Flow 0.918 ± 0.003 0.873 ± 0.004 453 233
Table 17: Ablation study on PhysioNet Sepsis dataset
We followed experimental protocol suggested by Kidger et al. (2020), and its Github repository [^17] https://github.com/patrick-kidger/NeuralCDE. For the benchmark methods, we used the reported performance in Jhin et al. (2022). Table 17 presents a comprehensive ablation study comparing various flow configurations integrated with Neural CDEs. We train for $200$ epochs with a batch size of $1024$ and a learning rate of $0.001$. The hyperparameters are optimized by grid search in the number of layers in vector field $n_{l}\in\left\{1,2,3,4\right\}$ and hidden vector dimensions $n_{h}\in\left\{16,32,64,128\right\}$. The best hyperparameters are highlighted with bold font in Table 21 and 21. Here it is acknowledged that increasing $n_{h}$ magnify memory and complexity instead of $n_{l}$. However, a more complex model does not necessarily assure superior performance.
Appendix E Detailed Results for “Forecasting with the Partial Observations”
We used two datasets for the forecasting with partial observations scenario. For the MuJoCo dataset, we followed experimental protocol suggested by in Jhin et al. (2023) and Jhin et al. (2024). We used reported performance of benchmark methods and source code from the Github repositories [^18] https://github.com/alflsowl12/LEAP. For the Google dataset, We followed experimental protocol and reported performance established by Jhin et al. (2022). Source code from its Github repository [^19] https://github.com/sheoyon-jhin/EXIT was utilized to conduct experiment. We conducted hyperparameter tuning for regular data and the same set of hyperparameters was applied to the scenarios entailing 30%, 50%, and 70% missing data.
MuJoCo.
The Multi-Joint dynamics with Contact (MuJoCo) dataset, developed by Todorov, Erez, and Tassa (2012), is based on the Hopper configuration in the DeepMind control suite (Tassa et al. 2018). This dataset comprises 10,000 simulations, each represented as a 14-dimensional time series with 100 data points sampled at regular intervals. In our experimental protocol, the initial 50 data points in each time series were used to predict the subsequent 10 data points. To introduce complexity and simulate diverse conditions, we systematically excluded a proportion of the data points—30%, 50%, and 70%—from each time series, as detailed in Jhin et al. (2023) and Jhin et al. (2024).
Configuration Test MSE Memory (MB)
Regular 30% Dropped 50% Dropped 70% Dropped
Conventional MLP 0.008 ± 0.002 0.009 ± 0.001 0.009 ± 0.001 0.009 ± 0.001 138
ResNet Flow 0.008 ± 0.003 0.008 ± 0.001 0.008 ± 0.001 0.008 ± 0.001 198
GRU Flow 0.006 ± 0.001 0.008 ± 0.001 0.008 ± 0.001 0.008 ± 0.001 318
Coupling Flow 0.008 ± 0.001 0.008 ± 0.001 0.008 ± 0.001 0.008 ± 0.001 409
Table 18: Forecasting performance and memory usage on MuJoCo dataset
Table 18 summarize the ablation study of the proposed method on MuJoCo dataset. In the proposed method, the training spanned $500$ epochs, with a batch size configured at $1024$ and a learning rate set to $0.001$. The process of hyperparameter optimization was conducted through grid search, concentrating on the number of layers $n_{l}\in\left\{1,2,3,4\right\}$, and the dimensions of the hidden vector $n_{h}\in\left\{16,32,64,128\right\}$. The hyperparameters that were identified as optimal are conspicuously presented in bold within Table 23. The same set of hyperparameters was applied to the irregularly-sampled scenarios.
Google.
The Google stock data (Jhin et al. 2022) includes transaction volumes of Google, along with metrics such as high, low, open, close, and adjusted closing prices. This dataset spans from 2011 to 2021, and the objective is to utilize historical time series data from the preceding 50 days to forecast the high, low, open, close, and adjusted closing prices for the next 10-day period. We followed the experimental protocol and reported performances of benchmark methods from Jhin et al. (2022). Similarly to the previous experiment, we excluded partial observations in data —30%, 50%, and 70%— in order to make more challenging situations.
Configuration Test MSE Memory (MB)
Regular 30% Dropped 50% Dropped 70% Dropped
Conventional MLP 0.0012 ± 0.0001 0.0011 ± 0.0001 0.0011 ± 0.0001 0.0014 ± 0.0001 52
ResNet Flow 0.0011 ± 0.0001 0.0010 ± 0.0001 0.0011 ± 0.0001 0.0013 ± 0.0001 62
GRU Flow 0.0012 ± 0.0001 0.0010 ± 0.0001 0.0011 ± 0.0001 0.0013 ± 0.0001 81
Coupling Flow 0.0010 ± 0.0001 0.0010 ± 0.0001 0.0009 ± 0.0001 0.0013 ± 0.0002 119
Table 19: Forecasting performance and memory usage on Google dataset
Table 19 summarize the ablation study of the proposed method on Google dataset. In the proposed method, the training spanned $200$ epochs, with a batch size configured at $64$ and a learning rate set to $0.001$. The process of hyperparameter optimization was conducted through grid search, concentrating on the number of layers $n_{l}\in\left\{1,2,3,4\right\}$, and the dimensions of the hidden vector $n_{h}\in\left\{16,32,64,128\right\}$. Through grid search, the optimal hyperparameters are highlighted in bold within Table 23. The same set of hyperparameters was applied to the irregularly-sampled scenarios.
$n_{l}$ $n_{h}$ Conventional MLP ResNet Flow GRU Flow Coupling Flow
Test AUROC Memory # Params Test AUROC Memory # Params Test AUROC Memory # Params Test AUROC Memory # Params
1 16 0.803 ± 0.018 225 23.81 0.867 ± 0.008 225 23.84 0.870 ± 0.014 225 24.15 0.887 ± 0.005 225 24.13
32 0.851 ± 0.004 227 83.46 0.878 ± 0.003 227 83.52 0.878 ± 0.003 252 84.64 0.880 ± 0.006 286 84.61
64 0.872 ± 0.004 242 313.35 0.889 ± 0.001 270 313.47 0.887 ± 0.005 355 317.76 0.883 ± 0.004 424 317.70
128 0.872 ± 0.002 345 1215.49 0.889 ± 0.006 401 1215.75 0.885 ± 0.004 571 1232.51 0.887 ± 0.005 710 1232.39
2 16 0.870 ± 0.010 225 24.35 0.900 ± 0.005 225 24.39 0.895 ± 0.003 225 24.42 0.893 ± 0.003 225 24.67
32 0.894 ± 0.003 227 85.57 0.901 ± 0.003 227 85.63 0.898 ± 0.003 252 85.70 0.903 ± 0.006 293 86.72
64 0.897 ± 0.005 256 321.67 0.905 ± 0.006 284 321.79 0.906 ± 0.003 355 321.92 0.900 ± 0.004 439 326.02
128 0.881 ± 0.005 373 1248.51 0.894 ± 0.002 430 1248.77 0.896 ± 0.005 571 1249.03 0.895 ± 0.005 738 1265.41
3 16 0.892 ± 0.004 225 24.90 0.905 ± 0.006 225 24.93 0.905 ± 0.004 225 24.69 0.905 ± 0.003 225 25.22
32 0.904 ± 0.005 227 87.68 0.912 ± 0.003 227 87.75 0.913 ± 0.004 252 86.75 0.916 ± 0.002 300 88.83
64 0.907 ± 0.005 270 329.99 0.912 ± 0.002 299 330.11 0.913 ± 0.002 355 326.08 0.918 ± 0.003 453 334.34
128 0.881 ± 0.002 401 1281.54 0.912 ± 0.004 458 1281.79 0.908 ± 0.002 570 1265.54 0.903 ± 0.006 766 1298.43
4 16 0.877 ± 0.007 225 25.44 0.906 ± 0.008 225 25.47 0.907 ± 0.004 225 24.96 0.911 ± 0.005 229 25.76
32 0.907 ± 0.003 227 89.79 0.915 ± 0.003 230 89.86 0.914 ± 0.002 252 87.81 0.915 ± 0.003 308 90.95
64 0.912 ± 0.011 284 338.31 0.910 ± 0.004 313 338.43 0.916 ± 0.003 355 330.24 0.911 ± 0.004 467 342.66
128 0.891 ± 0.003 429 1314.56 0.909 ± 0.005 487 1314.82 0.912 ± 0.004 571 1282.05 0.904 ± 0.005 795 1331.46
Table 20: Results of hyperparameter tuning for PhysioNet Sepsis Dataset with OI (Memory usage in MB, and the number of parameter in ×$\times 10^{3}$ unit.)
$n_{l}$ $n_{h}$ Conventional MLP ResNet Flow GRU Flow Coupling Flow
Test AUROC Memory # Params Test AUROC Memory # Params Test AUROC Memory # Params Test AUROC Memory # Params
1 16 0.777 ± 0.018 114 13.76 0.826 ± 0.007 114 13.79 0.827 ± 0.007 128 14.10 0.838 ± 0.008 145 14.08
32 0.815 ± 0.007 121 46.21 0.838 ± 0.008 135 46.27 0.845 ± 0.006 178 47.39 0.841 ± 0.007 212 47.36
64 0.815 ± 0.003 165 169.47 0.828 ± 0.004 193 169.60 0.841 ± 0.006 278 173.89 0.833 ± 0.003 348 173.83
128 0.790 ± 0.018 259 649.47 0.834 ± 0.004 316 649.73 0.825 ± 0.008 485 666.50 0.837 ± 0.010 624 666.37
2 16 0.835 ± 0.005 114 14.31 0.861 ± 0.005 114 14.34 0.860 ± 0.004 128 14.37 0.861 ± 0.004 148 14.63
32 0.839 ± 0.008 128 48.32 0.860 ± 0.004 142 48.39 0.862 ± 0.006 178 48.45 0.866 ± 0.001 219 49.47
64 0.830 ± 0.004 179 177.79 0.856 ± 0.005 208 177.92 0.851 ± 0.006 278 178.05 0.861 ± 0.003 362 182.15
128 0.820 ± 0.007 287 682.50 0.833 ± 0.008 344 682.75 0.848 ± 0.013 486 683.01 0.845 ± 0.008 653 699.39
3 16 0.824 ± 0.018 114 14.85 0.860 ± 0.008 114 14.88 0.865 ± 0.011 128 14.64 0.867 ± 0.007 152 15.17
32 0.861 ± 0.016 135 50.43 0.871 ± 0.006 149 50.50 0.875 ± 0.004 178 49.51 0.869 ± 0.004 226 51.59
64 0.844 ± 0.006 193 186.11 0.859 ± 0.010 222 186.24 0.869 ± 0.003 278 182.21 0.865 ± 0.008 376 190.47
128 0.818 ± 0.013 316 715.52 0.852 ± 0.006 373 715.78 0.858 ± 0.008 486 699.52 0.856 ± 0.010 681 732.42
4 16 0.799 ± 0.070 114 15.39 0.853 ± 0.010 117 15.43 0.854 ± 0.006 128 14.91 0.871 ± 0.009 155 15.71
32 0.848 ± 0.017 142 52.55 0.875 ± 0.006 156 52.61 0.867 ± 0.007 178 50.56 0.873 ± 0.004 233 53.70
64 0.845 ± 0.015 208 194.43 0.863 ± 0.006 236 194.56 0.863 ± 0.006 278 186.37 0.864 ± 0.005 390 198.79
128 0.823 ± 0.004 345 748.55 0.846 ± 0.005 402 748.80 0.858 ± 0.011 486 716.03 0.858 ± 0.010 710 765.44
Table 21: Results of hyperparameter tuning for PhysioNet Sepsis Dataset without OI (Memory usage in MB, and the number of parameter in ×$\times 10^{3}$ unit.)
$n_{l}$ $n_{h}$ Conventional MLP ResNet Flow GRU Flow Coupling Flow
Test MSE Memory # Params Test MSE Memory # Params Test MSE Memory # Params Test MSE Memory # Params
1 16 0.020 ± 0.001 36 5.10 0.018 ± 0.001 42 5.13 0.017 ± 0.002 57 5.44 0.016 ± 0.002 69 5.42
32 0.012 ± 0.001 49 18.89 0.012 ± 0.003 62 18.96 0.009 ± 0.001 92 20.08 0.011 ± 0.000 118 20.05
64 0.009 ± 0.002 77 72.59 0.009 ± 0.003 105 72.72 0.006 ± 0.001 165 77.01 0.008 ± 0.000 214 76.94
128 0.008 ± 0.001 138 284.43 0.008 ± 0.002 198 284.69 0.006 ± 0.000 318 301.45 0.008 ± 0.002 409 301.33
2 16 0.022 ± 0.002 38 5.65 0.020 ± 0.001 45 5.68 0.017 ± 0.003 57 5.71 0.019 ± 0.002 72 5.97
32 0.016 ± 0.003 54 21.01 0.013 ± 0.002 67 21.07 0.010 ± 0.001 92 21.13 0.012 ± 0.002 123 22.16
64 0.012 ± 0.001 87 80.91 0.010 ± 0.002 114 81.04 0.010 ± 0.001 165 81.17 0.011 ± 0.001 225 85.26
128 0.011 ± 0.001 158 317.45 0.010 ± 0.001 219 317.71 0.008 ± 0.001 319 317.97 0.010 ± 0.000 430 334.35
3 16 0.039 ± 0.021 41 6.19 0.021 ± 0.005 47 6.22 0.016 ± 0.003 57 5.98 0.016 ± 0.004 74 6.51
32 0.033 ± 0.026 59 23.12 0.014 ± 0.002 72 23.18 0.011 ± 0.002 92 22.19 0.012 ± 0.001 128 24.27
64 0.032 ± 0.027 97 89.23 0.011 ± 0.000 125 89.36 0.010 ± 0.002 165 85.33 0.010 ± 0.001 234 93.58
128 0.013 ± 0.004 179 350.48 0.011 ± 0.002 239 350.73 0.008 ± 0.000 319 334.48 0.011 ± 0.001 450 367.37
4 16 0.042 ± 0.018 43 6.73 0.024 ± 0.004 50 6.77 0.022 ± 0.009 57 6.25 0.022 ± 0.006 77 7.05
32 0.034 ± 0.026 64 25.23 0.014 ± 0.001 78 25.29 0.014 ± 0.003 92 23.25 0.012 ± 0.003 132 26.38
64 0.032 ± 0.027 107 97.55 0.012 ± 0.002 134 97.68 0.011 ± 0.002 165 89.49 0.009 ± 0.002 245 101.90
128 0.031 ± 0.028 199 383.50 0.011 ± 0.002 260 383.76 0.008 ± 0.001 319 350.99 0.010 ± 0.000 471 400.40
Table 22: Results of hyperparameter tuning for MuJoCo Dataset (Memory usage in MB, and the number of parameter in ×$\times 10^{3}$ unit.)
$n_{l}$ $n_{h}$ Conventional MLP ResNet Flow GRU Flow Coupling Flow
Test MSE Memory # Params Test MSE Memory # Params Test MSE Memory # Params Test MSE Memory # Params
1 16 0.0017 ± 0.0001 6 2.97 0.0019 ± 0.0005 8 3.00 0.0017 ± 0.0004 12 3.30 0.0012 ± 0.0000 15 3.29
32 0.0014 ± 0.0000 10 11.05 0.0013 ± 0.0001 13 11.11 0.0014 ± 0.0001 21 12.23 0.0011 ± 0.0000 26 12.20
64 0.0013 ± 0.0000 17 42.57 0.0012 ± 0.0000 24 42.70 0.0012 ± 0.0000 40 46.98 0.0011 ± 0.0001 51 46.92
128 0.0013 ± 0.0000 35 167.05 0.0012 ± 0.0001 51 167.30 0.0012 ± 0.0000 81 184.07 0.0011 ± 0.0001 102 183.94
2 16 0.0015 ± 0.0001 7 3.51 0.0022 ± 0.0009 8 3.54 0.0017 ± 0.0004 12 3.58 0.0011 ± 0.0001 15 3.83
32 0.0014 ± 0.0000 11 13.16 0.0014 ± 0.0003 14 13.22 0.0014 ± 0.0001 21 13.29 0.0011 ± 0.0000 28 14.31
64 0.0013 ± 0.0000 19 50.89 0.0012 ± 0.0001 26 51.02 0.0012 ± 0.0001 40 51.14 0.0011 ± 0.0000 54 55.24
128 0.0013 ± 0.0000 41 200.07 0.0011 ± 0.0000 56 200.33 0.0012 ± 0.0000 81 200.58 0.0011 ± 0.0000 108 216.97
3 16 0.0018 ± 0.0003 7 4.06 0.0017 ± 0.0007 9 4.09 0.0015 ± 0.0004 12 3.85 0.0011 ± 0.0001 16 4.38
32 0.0014 ± 0.0001 12 15.27 0.0013 ± 0.0003 16 15.34 0.0014 ± 0.0002 21 14.34 0.0011 ± 0.0000 29 16.42
64 0.0013 ± 0.0000 22 59.21 0.0013 ± 0.0001 29 59.34 0.0012 ± 0.0001 40 55.30 0.0011 ± 0.0000 57 63.56
128 0.0012 ± 0.0000 47 233.10 0.0011 ± 0.0001 62 233.35 0.0012 ± 0.0000 82 217.10 0.0011 ± 0.0001 113 249.99
4 16 0.0015 ± 0.0001 8 4.60 0.0022 ± 0.0013 10 4.63 0.0018 ± 0.0006 12 4.12 0.0011 ± 0.0001 16 4.92
32 0.0013 ± 0.0000 13 17.38 0.0016 ± 0.0005 17 17.45 0.0012 ± 0.0001 21 15.40 0.0011 ± 0.0000 30 18.54
64 0.0013 ± 0.0000 25 67.53 0.0012 ± 0.0001 32 67.66 0.0012 ± 0.0001 40 59.46 0.0011 ± 0.0000 59 71.88
128 0.0012 ± 0.0000 52 266.12 0.0011 ± 0.0000 68 266.38 0.0012 ± 0.0001 82 233.61 0.0010 ± 0.0000 119 283.02
Table 23: Results of hyperparameter tuning for Google Dataset (Memory usage in MB, and the number of parameter in ×$\times 10^{3}$ unit.)